### v1

# AC-LIO: Towards Asymptotic and Consistent Convergence in LiDAR-Inertial Odometry

Abstract
Existing LiDAR-Inertial Odometry (LIO) frameworks typically utilize prior state trajectories derived from IMU integration to compensate for the motion distortion within LiDAR frames, and demonstrate outstanding accuracy and stability in regular low-speed and smooth scenes. However, in high-speed or intense motion scenarios, the residual distortion may increase due to the limitation of IMU’s accuracy and frequency, which will degrade the consistency between the LiDAR frame with its represented geometric environment, leading pointcloud registration to fall into local optima and consequently increasing the drift in long-time and large-scale localization.
To address the issue, we propose a novel asymptotically and consistently converging LIO framework called AC-LIO. First, during the iterative state estimation, we backwards propagate the update term based on the prior state chain, and asymptotically compensate the residual distortion before next iteration. Second, considering the weak correlation between the initial error and motion distortion of current frame, we propose a convergence criteria based on pointcloud constraints to control the back propagation. The approach of guiding the asymptotic distortion compensation based on convergence criteria can promote the consistent convergence of pointcloud registration and increase the accuracy and robustness of LIO. Experiments show that our AC-LIO framework, compared to other state-of-the-art frameworks, effectively promotes consistent convergence in state estimation and further improves the accuracy of long-time and large-scale localization and mapping.
I INTRODUCTION
Figure 1: The basic principle of AC-LIO framework. Conventional LIO typically consider initial distortion compensation via IMU integration, yet the residual distortion may prevent the registration from further convergence. In contrast, our AC-LIO employs an asymptotic backpropagation strategy to promote consistent convergence in state estimation.
With the development of Simultaneous Localization and Mapping (SLAM) technology, the perception of robots on the external environment and their own motion is becoming more and more precise [1, 2]. As a key branch, LiDAR-Inertial Odometry (LIO) has gained widespread application in autonomous driving, unmanned logistics and other fields, for its high robustness and accuracy [3, 4]. Compared to Visual-Inertial Odometry (VIO), the LIO system can directly obtain depth observations of surroundings [5], without the need to estimate the depth of feature points [6, 7], and exhibits better adaptability to different lighting conditions [8, 9].
However, as LIO system continues to evolve, there still remain several fundamental issues to be addressed [10, 11], one of which is the conflict between continuous motion and rigid pointcloud registration [12]. Since the acquisition of LiDAR stream is often accompanied with the ego-motion of vehicle, the laser points with different timestamps actually lie in different coordinate frames.
Existing frameworks typically only consider using IMU integration for the initial compensation of motion distortion [13, 14, 15], without further optimization of the residual distortion in state estimation. But in high-speed or violent motion scenarios, the residual motion distortion tends to increase due to the limitations of IMU’s accuracy and frequency, which would suppress further convergence of pointcloud registration and decrease odometry accuracy in long-term, large-scale scenarios [16].
To address this, we propose AC-LIO, a novel asymptotically and consistently converging LIO framework. First, during the iterative state estimation, we backwards propagate the current update term based on the prior state chain, and asymptotically compensate the residual distortion before next iteration. This approach could enhance the consistency between the current LiDAR frame and the represented geometric environment, and avoid the pointcloud registration from falling into local optimum due to the residual distortion, thereby promoting the consistent convergence of state estimation and improving the accuracy of long-time and large-scene localization. Compared to other strategies such as continuous-time estimation or frame segmentation, our method can be seamlessly embedded into the iterative Kalman Filter framework [17], offering great flexibility and scalability.
Second, considering the weak correlation between the initial error and motion distortion of current frame, we also propose a convergence criteria based on the point-plane residuals to control the back propagation. This strategy ensures, We use this criteria to evaluate the convergence of the pointcloud registration, ensuring the backpropagation and asymptotic compensation only happens under the condition of sufficient convergence in previous LiDAR frame, and thus guaranteeing the robustness of LIO system.
In summary, our main contributions are as follows:
  • We implement a novel asymptotically and consistently converging LIO framework. During the iterative update, we back propagate the current update term based on the prior state chain, and asymptotically compensate the residual distortion before next iteration, promoting the consistent convergence of registration and improving the accuracy in long-time and large-scale scenarios.
  • We introduce a convergence criteria based on point-plane residuals to control the back propagation. We utilize the criteria to evaluate the convergence of registration, ensuring the back propagation to be performed only upon sufficient convergence of the previous registration, thus further enhancing the system robustness.
  • Our method can be easily and seamlessly integrated into traditional iterative Kalman Filter framework. Compared to other optimization strategies for motion distortion, our framework introduce no extra estimates or states, making the asymptotic compensation less computationally expensive but highly flexible and scalable.
  • We conduct extensive experiments on AC-LIO in a series of real-world datasets with various scenarios. The results show that our proposed AC-LIO exhibits higher accuracy in long-term localization compared to the state-of-the-art open-source methods.
II RELATED WORK
Since the motion distortion can degrade the consistency between the pointcloud and the geometric environment, and thus suppress the sufficient convergence of registration. To mitigate the adverse effect of motion distortion, existing LiDAR(-inertial) odometry adopt various strategies to appropriately compensate for this distortion.
For LiDAR-only odometry, most studies utilize the constant velocity model to compensate for motion aberrations. Studies such as LOAM [18] and Lego-LOAM [19] have achieved significant effects with the assumption of constant velocity motion during the corresponding time of LiDAR frames. However, this assumption only works effectively when LiDAR frame interval is enough short and the motion is slow and smooth. In case of drastic velocity variation, the constant velocity model cannot accurately characterize the current motion state, and this will result in increased drift and even odometry divergence.
As for LiDAR-inertial odometry, the IMU observation can provide relatively reliable prior state information. This provides a new scheme for compensating motion distortion via IMU integration. FAST-LIO [20] employs the IMU state backpropagation to correct the distorted pointcloud, and then conducts state estimation through Error State Kalman Filter. LIO-SAM [21] utilizes the IMU pre-integration to de-skew pointcloud, and combines multiple constraints including LiDAR, IMU, and GPS through the factor graph. Although these methods can mitigate the impact of pointcloud distortion to some extent, there are still some limitations without considering the further optimization of residual motion distortion.
To characterize trajectorie more precisely in state estimation, many frameworks come up with new optimization solutions. As pure LiDAR framework, I2EKF-LO [22] utilizes additional iteration to specifically estimate motion distortion, this approach can suppress distortion but the extra iteration may sacrifice computational efficiency. CLINS [23] introduces the B-spline model [24] for continuous time state estimation. However, the unconstrained trajectory would not always conform to the B-spline model due to time resolution limitations in high-speed, dynamic motion. Moreover, the continuous-time optimization strategy may introduce a higher computational cost. Point-LIO [25] adopts a point-by-point update strategy to avoid the motion distortion issue and realize extremely high-frequency odometry. Compared to frame based registration, the point-by-point strategy is more likely to degenerate in adverse scenarios. FR-LIO [26] divides the LiDAR frame into several sub-frames based on IMU observation. This approach can increase the system update frequency, thus reducing the impact of motion distortion. Besides, FR-LIO strengthens the constraints between subframes via RTS smoother [27], but still faces the risk of subframe degradation.
In contrast, our proposed AC-LIO framework utilizes the convergence criteria to guide the backpropagation process, to achieve the asymptotic compensation of residual motion distortion. As a result, our approach can promote the consistent convergence of pointcloud registration during iterative update, further improving the accuracy over long-term and large-scale scenarios while ensuring the system robustness.
III METHODOLOGY
III-A System Overview
Figure 2: The system overview of our AC-LIO framework. During iterative ESKF, asymptotic compensation of pointcloud distortion will be conducted based on the convergence criteria, to promote consistent convergence of state estimation.
The system overview of our AC-LIO framework is shown in Fig. 2. IMU observations are fed into the forward propagation module for integration to obtain prior trajectories. Then, the LiDAR data utilize the prior state chain for initial compensation, next passed into the Error State Kalman Filter (ESKF) for state estimation [28], where the average point-to-plane residual (APR) is calculated after each update.
During the iteration, if the APR of previous LiDAR frame is lower than the convergence criteria and the current APR exceeds it, then the backpropagation of update residual will be executed. After backpropagation, the LiDAR frame needs to be re-compensated to promote the consistent convergence of state estimation. Otherwise, if the conditions are not met, the backpropagation will be skipped and iteration continues. Once reaching the maximum of iterations, the system will merge the LiDAR frame under the posterior pose into pointcloud map and publish the current odometry.
III-B ESKF State Estimation
For efficiency and accuracy concerns, the AC-LIO we proposed utilizes the ESKF framework for state estimation. Based on the kinematic model and IMU measurement model, the state $\mathbf{x}$ , input $\mathbf{i}$, and noise $\mathbf{n}$ can be defined as follows:
$\begin{split}\mathbf{x}&\doteq\begin{bmatrix}\mathbf{R}_{I}^{G}&\mathbf{p}_{I}^{G}&\mathbf{v}_{I}&\mathbf{b}_{\bm{\omega}}&\mathbf{b}_{\mathbf{a}}&\mathbf{g}^{G}\end{bmatrix}^{T}\\ \mathbf{i}&\doteq\begin{bmatrix}{\bm{\omega}}_{m}&{\mathbf{a}}_{m}\end{bmatrix}^{T}\\ \mathbf{n}&\doteq\begin{bmatrix}\mathbf{n}_{\bm{\omega}}&\mathbf{n}_{\mathbf{a}}&\mathbf{n}_{\mathbf{b}\bm{\omega}}&\mathbf{n}_{\mathbf{b}\mathbf{a}}\end{bmatrix}^{T}\end{split}$ (1)
Where $\mathit{G}$ and $\mathit{I}$ denote the global and the IMU frames, $\mathbf{R}_{I}^{G}$, $\mathbf{p}_{I}^{G}$ and $\mathbf{v}_{I}^{G}$ are IMU’s orientation, position and linear velocity in $\mathit{G}$, and $\mathbf{g}^{G}$ is the local gravity. The $\mathbf{n}_{\bm{\omega}}$, $\mathbf{n}_{\mathbf{a}}$ are noises of angular velocity ${\bm{\omega}}_{m}$ and linear acceleration ${\mathbf{a}}_{m}$, the $\mathbf{b}_{\bm{\omega}}$, $\mathbf{b}_{\mathbf{a}}$ are IMU bias with $\mathbf{n}_{\mathbf{b}\bm{\omega}}$ and $\mathbf{n}_{\mathbf{b}\mathbf{a}}$ as their noises in Gaussian-Markov model.
[…]
$\begin{gathered}\quad\quad\delta\mathbf{x}_{i}\doteq\begin{bmatrix}\delta{\bm{\theta}}&\delta\mathbf{p}_{I}^{G}&\delta\mathbf{v}_{I}&\delta\mathbf{b}_{\bm{\omega}}&\delta\mathbf{b}_{\mathbf{a}}&\delta\mathbf{g}^{G}\end{bmatrix}^{T}\\ \quad\delta\hat{\mathbf{x}}_{i+1}=\mathbf{F_{a}}\delta\check{\mathbf{x}}_{i}+\mathbf{F_{n}}\mathbf{n}_{i}\\ \quad\hat{\mathbf{P}}_{i+1}=\mathbf{F_{a}}\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}+\mathbf{F_{n}}\mathbf{Q}\mathbf{F_{n}}^{T}\end{gathered}$ (2)
Denote $\mathbf{T}_{L}^{I}$ is the extrinsic between LiDAR and IMU, then the pointcloud can be transformed to frame $\mathit{I}$ via the prior pose chain. For a point $\mathbf{p}$ in LiDAR frame, its neighboring cluster can be searched by KNN in pointcloud map. A plane is then fitted to this cluster and we obtain the normal vector $\mathbf{u}$. Note one point of the cluster as $\mathbf{q}$ and $(\cdot)_{\wedge}$ as the upper triangular matrix of vector, then the point-plane residual $\mathbf{z}$ and measurement matrix $\mathbf{H}$ are as follows:
$\begin{gathered}\mathbf{z}=\mathbf{u}(\hat{\mathbf{p}}^{G}-\mathbf{q}^{G})^{T}\\ \mathbf{H}=\begin{bmatrix}-\mathbf{u}\hat{\mathbf{R}}_{I}^{G}(\mathbf{T}_{L}^{I}\mathbf{p}^{L})_{\wedge}&\mathbf{u}&\mathbf{0}_{3\times 12}\end{bmatrix}\end{gathered}$ (3)
Considering the nonlinear errors in the $\mathbf{F_{a}}$, $\mathbf{F_{n}}$ and $\mathbf{H}$, the estimation may require iteration to promote convergence. Let $\mathbf{A}$ denote the Jacobian of update component with respect to $\delta\mathbf{x}$, and according to [13], the update equation is:
$\begin{gathered}\mathbf{K}=\mathbf{P}\mathbf{H}^{T}(\mathbf{H}\mathbf{P}\mathbf{H}^{T}+\mathbf{R})^{-1}\\ \hat{\mathbf{x}}_{i+1}^{j+1}=\hat{\mathbf{x}}_{i+1}^{j}\boxplus[-\mathbf{K}\mathbf{z}-(\mathbf{I}-\mathbf{K}\mathbf{H})(\mathbf{A}^{j})^{-1}(\hat{\mathbf{x}}_{i+1}^{j}\boxminus\hat{\mathbf{x}}_{i+1})]\end{gathered}$ (4)
Once the update term is less than a preset threshold, the solution can be considered to converge.
$\begin{gathered}\check{\mathbf{x}}_{i+1}\Leftarrow\hat{\mathbf{x}}_{i+1}^{j+1},\;\check{\mathbf{P}}_{i+1}\Leftarrow(\mathbf{I}-\mathbf{K}\mathbf{H})\hat{\mathbf{P}}_{i+1}\end{gathered}$ (5)
With the iterative update of ESKF, our AC-LIO could further promote the consistent convergence of the state estimation via the asymptotic distortion compensation and convergence criteria introduced in following sections.
III-C Asymptotic Distortion Compensation
In smooth motion scenarios, an initial compensation of motion distortion via IMU integration can basically satisfy the accuracy demand. However, when encountering violent motion scenarios, the residual motion distortion may increase due to the limitation of IMU’s frequency and accuracy. Compared to the noise of LiDAR observations, motion distortion may introduce bias in pointcloud registration, making it difficult to utilize the statistical properties of pointcloud to mitigate the adverse effects on state estimation.
With reference to the fixed interval smooth strategy, we selectively back propagate the update term $\delta\mathbf{x}$ based on the prior state chain during iteration. The noise between adjacent states can be expressed as follows:
$\begin{gathered}\delta\check{\mathbf{x}}_{i}\sim\mathcal{N}(\mathbf{0}_{18\times 1},\check{\mathbf{P}}_{i})\\ \delta\hat{\mathbf{x}}_{i+1}\sim\mathcal{N}(\mathbf{0}_{18\times 1},\mathbf{F_{a}}\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}+\mathbf{F_{n}}\mathbf{Q}\mathbf{F_{n}}^{T})\end{gathered}$ (6)
Figure 3: The on-manifold propagation. The noise of current state involves the process noise and the propagation of previous state, the update term can be recursive backward according to this relationship.
During the on-manifold propagation of state $\mathbf{x}$, the noise of current error state $\delta\hat{\mathbf{x}}_{i+1}$ consists of the propagation of previous state $\delta\hat{\mathbf{x}}_{i}$ and the accumulated process noise. According to RTS smooth strategy [27], the contribution of previous state in current update term just corresponds to the proportion of the propagation covariance $\mathbf{F_{a}}\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}$ within current covariance $\hat{\mathbf{P}}_{i+1}$. Consequently, the gain matrix $\mathbf{G}$ and the backward recursion of $\delta\mathbf{x}$ can be denoted as follows:
$\begin{gathered}\mathbf{G}=\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}\hat{\mathbf{P}}_{i+1}^{-1}\\ \delta\hat{\mathbf{x}}_{i}^{j}=\mathbf{G}\cdot\delta\hat{\mathbf{x}}_{i+1}^{j}\end{gathered}$ (7)
Furthermore, since there may exist numerous IMU states to maintain within the timeperiod of LiDAR frame, a sequential backward recursion can characterize the smoothed trajectory precisely. However, from an overall perspective, this treatment will increase much computational cost in iterative ESKF, but cannot bring additional accuracy improvement in subsequent asymptotic distortion compensation. Compared to state-by-state backpropagation, selectively updating several anchors in the prior state chain can further reduce the computational cost with little accuracy loss. To this end, we extend the backwards recursion to any states within the LiDAR frame, thereby reducing the computational complexity of backpropagation and distortion compensation.
$\begin{gathered}\mathbf{F_{a}}^{i|i+n}=\prod_{k=i}^{i+n-1}\mathbf{F_{a}}^{k|k+1}\\ \mathbf{G}_{i}^{i+n}=\check{\mathbf{P}}_{i}(\mathbf{F_{a}}^{i|i+n})^{T}(\hat{\mathbf{P}}_{i+n})^{-1}\\ \check{\mathbf{x}}_{i}^{j}=\check{\mathbf{x}}_{i}\boxplus(\mathbf{G}_{i}^{i+n}\cdot\delta\hat{\mathbf{x}}_{i+n}^{j})\end{gathered}$ (8)
After each backpropagation of the update term $\hat{\mathbf{x}}_{i+n}^{j}$, the residual pointcloud distortion can continue to be compensated based on the smoothed trajectory. For the point $\mathbf{p}^{L}$ in LiDAR frame, and the state corresponding to its timestamp $t$ after the $j$th backpropagation can be denoted as $\hat{\mathbf{x}}_{t}^{j}$, then the updated coordinates can be noted as:
$\begin{gathered}\hat{\mathbf{p}}_{j}^{G}=\mathbf{T}(\hat{\mathbf{x}}_{t}^{j})\mathbf{T}_{L}^{I}\mathbf{p}^{L}\end{gathered}$ (9)
Where $\mathbf{T}(\mathbf{x})$ represents the transformation matrix of state $\mathbf{x}$ in frame $G$. As the iterative ESKF converges, the final solution of the backward propagation can be expressed as:
$\begin{gathered}\bar{\mathbf{x}}_{i}\Leftarrow\check{\mathbf{x}}_{i}^{j+1},\;\bar{\mathbf{P}}_{i}\Leftarrow\mathbf{G}_{i}^{i+n}\check{\mathbf{P}}_{i+n}(\mathbf{G}_{i}^{i+n})^{T}\end{gathered}$ (10)
In practice, if there is no need to reuse the previous state queue in next frame, the covariance update after backpropagation can be skipped for efficiency.
III-D Convergence Criteria
In order to exploit the statistical properties of the pointcloud, during the iterative state estimation, we select the mean value of the point-to-plane residual (abbreviated as APR above), as the criteria for evaluating the convergence of pointcloud registration. For simplification, assuming the pointcloud registration fully converges to the global optimum, considering the ranging accuracy of LiDAR and the construction of point-to-plane residual, we can roughly model the expectation of the APR, to provide a threshold reference for the convergence criteria.
Denote the fitted plane in pointcloud map as $S$ with its normal vector as $\mathbf{u}$, and the point-plane residual as $\mathbf{z}$. Then, the incident angle $\phi$ of point $\mathbf{p}$ with respect to plane $S$ can be expressed as:
$\begin{gathered}\phi=\arccos\dfrac{|\mathbf{u}\cdot\mathbf{p}|}{|\mathbf{u}|\cdot|\mathbf{p}|}\end{gathered}$ (11)
When the laser beam is perpendicular to plane $S$, then the incident angle $\phi$ will be zero. Assuming that the ranging error $\mathbf{n_{p}}$ of laser point follows Gaussian distribution with zero mean and $\sigma$ variance, then the residual $\mathbf{z}_{\perp}$ will follow chi-squared distribution and can be expressed as follows:
$\begin{gathered}\mathbf{p}^{L}=\mathbf{p}_{gt}^{L}+\mathbf{n_{p}},\;\mathbf{n_{p}}\sim\mathcal{N}(0,\sigma)\\ \mathbf{z}_{\perp}=|\mathbf{p}_{gt}^{L}-\mathbf{p}^{L}|\\ \frac{\mathbf{z}_{\perp}^{2}}{\sigma^{2}}\sim\chi^{2}(1),\;\mathbb{E}(\mathbf{z}_{\perp})=\sigma\end{gathered}$ (12)
As the surrounding environment varies during ego-motion, the distribution of normal vector $\mathbf{n}$ of plane $S$ will also change. Eventually, the variation of the geometrical scene will be reflected in the distribution of the incident angle $\phi$. Assuming the incident angle is uniformly distributed over the domain $[0,\pi/2)$ , then the expected mean of the point-plane residual $\mathbf{z}$ can be expressed as:
$\begin{gathered}\mathbb{E}(\mathbf{z})=\frac{2}{\pi}\int_{0}^{\frac{\pi}{2}}\mathbb{E}(\mathbf{z}_{\perp})\cos{\phi}\;d\phi=\frac{2}{\pi}\sigma\end{gathered}$ (13)
Therefore, assuming the pointcloud registration converges to the global optimum, and only considering the LiDAR ranging error, the expectation of APR should be ${2\sigma}/{\pi}$. If the mean value of current point-plane residual is lower than this criteria, then the last registration can be considered as sufficiently converged.
According to the threshold of convergence criteria, we can easily determine the convergence degree of pointcloud registration, and control the backpropagation and the asymptotic distortion compensation. Noting that $\sigma_{t}=\mathbb{E}(\mathbf{z})$ and the selected threshold of convergence criteria as $\eta$, due to the factors such as dynamic objects and cumulative errors in pointcloud map, the threshold $\eta$ can be appropriately extended to the interval of $[\sigma_{t},2\sigma_{t}]$ in practice.
In particular, we perform two conditional judgments in iterative state estimation: First, evaluate the previous LiDAR frame, if the APR is less than the threshold, it is considered to be sufficiently converged and we can continue. Then, check the convergence of current registration, if the APR is still less than the threshold, there is no need to back-propagate for small residual distortion; otherwise, perform the backpropagation and asymptotic compensation, in order to promote the consistent convergence in state estimation.
The approach, on one hand, can ensure the backpropagation and asymptotic compensation only performed as the previous LiDAR frame converged well, thus enhancing the odometry robustness in long-time, large-scale scenario; on the other hand, can avoid lots of additional operations in unreliable or unnecessary conditions, further saving the system computational expense. The experiment results of our framework are shown in the next section.
IV EXPERIMENT RESULTS
In this section, we qualitatively and quantitatively evaluate AC-LIO framework through a series of experiments with various public and self-collected datasets. To examine the accuracy improvement of our approach in long-time and large-scale scenarios, and the effect of convergence criteria on system robustness, we conduct extensive experiments on the odometry accuracy and the validity of convergence criteria.
IV-A Odometry Accuracy Evaluation
In the odometry accuracy experiments, we compare AC-LIO framework with current state-of-the-art LIO systems including FAST-LIO2 and Point-LIO. All algorithms are implemented in C++ and employed on an i9 CPU desktop computer with the ROS operating system in Ubuntu. These methods only utilize CPU for computation without the assistance of GPU parallel computing. Besides, in the odometry accuracy experiments, we maintain the threshold of convergence criteria in the extended interval, i.e., $\eta\in[\sigma_{t},2\sigma_{t}]$. A brief introduction of the public and self-collected handheld datasets is provided below:
Figure 4: Our handheld sensor suite used in self-collected dataset, including the solid-state LiDAR Livox Avia with the built-in BMI088 IMU, and three other cameras.
TABLE I: Odometry Accuracy Evaluation - RMSE of ATE ($m$)
AC-LIO(Ours) FAST-LIO2 Point-LIO
Residence_5.2
0.415 0.505 1
Park_3.2
0.582 0.618
Subway_3.3
1.670 3.752
Mall_6.3
0.216 0.208 15.409
Garden_1005
0.773 1.053 0.917
Garden_1006
1.796 2.075 2.471
Garden_1008
1.152 1.577 1.590
  • 1
    –  means that the system severely diverged midway.
TABLE II: Odometry Accuracy Evaluation - End to end errors ($m$)
AC-LIO(Ours) FAST-LIO2 Point-LIO
Indoor_1
(~100m)
0.8929 1.1330 5.4006
Indoor_2
(~100m)
0.2267 0.2043 0.1636
Indoor_3
(~100m)
1.5959 3.0011 6.8438
Outdoor_1
(~300m)
0.1378 1.9007 0.7280
Outdoor_2
(~400m)
0.9840 2.1944 1.2607
Outdoor_3
(~500m)
2.6659 4.7023 4.6066
  • WHU-Helmet [30]: Collected on compact helmet system, the dataset contains complex large-scale scenes including wild, urban and tunnel, with challenges of drastic motion from wearer and repeated symmetrical geometric structures. The datasets used include Residence_5.2, Park_3.2, Subway_3.3, Mall_6.3.
  • BotanicGarden [31]: Mounted on all-terrain wheeled robot, this dataset consists of diverse natural scenes including riversides, narrow trails and bridges. The datasets used include the Garden series.
  • Handheld datasets: We collected multiple datasets of different scenarios and motion conditions, covering indoor and outdoor scenes like library, campus road and UAV airfield. The handheld data collection device is shown in Fig. 4. The datasets used include Indoor and Outdoor series.
In the evaluation of WHU-Helmet, Point-LIO diverged midway in the Residence_5.2 and Subway_3.3 due to severe motion combined with locally degraded scenarios, and diverged in Park_3.2 due to extensive occlusion by pedestrians. And in Mall_6.3, the system experienced significant drift during turns due to the reflection from glass walls of the shopping mall. In the first three datasets, our AC-LIO framework effectively resists the challenging scenarios mentioned above and outperformed the FAST-LIO in long-term, large-scale localization accuracy. Only in the last dataset of WHU-Helmet slightly lower than FAST-LIO.
For the BotanicGarden datasets, since the sensors are mounted on a wheeled robotic platform, the overall motion is smooth and suitable for evaluating the long-duration drift under regular motion conditions. Across the three LIO-rosbags provided by BotanicGarden, the AC-LIO demonstrates higher accuracy, exhibiting better convergence in long-term localization.
Next is the evaluation of handheld datasets. In Indoor_1 UAV airfield and Indoor_3 library scenarios, we conduct intense turning in narrow environment. And in the Indoor_2 garage scene, with smooth motion and highly structured surroundings, Point-LIO achieves higher closed-loop accuracy. Then in the Outdoor sequences, we handhold the device and attempt different motion conditions including running and sharp turns in campus road scenes.
Concrete experiment results are demonstrated in Table I, Table II and Fig. 5 . The results indicate the AC-LIO framework generally outperforms the FAST-LIO2 and Point-LIO in terms of accuracy during long-term, large-scale scenes, and exhibits high robustness under challenging scenarios.
Figure 5: Accuracy evaluation in campus road scene. The figure shows the result of Outdoor_1. We keep running after initialization to enhance the motion intensity and complete a closed loop about 300m. The closure trajectory and the map consistency suggest that our AC-LIO achieves better performance.
IV-B Convergence Criteria Validity Evaluation
To evaluate the validity of the proposed convergence criteria, we examine the odometry accuracy at different $\eta$ threshold levels. While keeping other parameters the same, we only adjust the threshold of convergence criteria. Due to the factors such as dynamic objects, cumulative pointcloud map errors, etc., we fine-tune the threshold $\eta$ within the interval of $[\sigma_{t},2\sigma_{t}]$. The specific experimental results are shown in Table III.
  • $\eta=0$: No backpropagation and asymptotic compensation is performed at this situation, reflecting the basic accuracy of the odometry.
  • $\eta=0.5\sigma_{t}$: Lower than the expected mean of pointcloud residuals, rarely perform backpropagation. (Only considering LiDAR ranging errors, $\mathbb{E}(\mathbf{z})=\sigma_{t}$)
  • $\eta\in[\sigma_{t},2\sigma_{t}]$: Recommended interval in consideration of other error factors, assuring the backpropagation and asymptotic compensation will be performed with sufficient convergence in previous frame.
  • $\eta=3\sigma_{t}$: Too high a threshold may lead to backpropagation in case of large initial errors, potentially damaging the system’s accuracy and robustness.
The experimental results show that, in most cases, selecting the threshold $\eta$ within the range of $[\sigma_{t},2\sigma_{t}]$ achieved higher accuracy. Overall, as the threshold increases from zero, the odometry accuracy initially improves because more sufficiently converged states are allowed to perform backpropagation and asymptotic compensation. However, as the threshold increases further, states with larger initial errors are also allowed to backpropagate, which reduces the effectiveness of the asymptotic compensation and negatively impacts odometry accuracy.
TABLE III: Convergence Criteria Validity Evaluation - RMSE of ATE ($m$)
00 0 $0.5\sigma_{t}$ $[\sigma_{t},2\sigma_{t}]$ $3\sigma_{t}$
Park_3.2
0.624 0.610 0.582 0.616
Subway_3.3
2.841 2.419 1.670 1.788
Garden_1006
1.901 1.709 1.796 1.892
Garden_1008
1.197 1.199 1.152 1.178
V CONCLUSIONS
In this paper, we propose AC-LIO, a novel asymptotically and consistently converging LIO framework. We reconsider the factors that may introduce errors in long-duration, large-scale scenarios, aiming to further enhance the convergence of the odometry. First, during iterative state estimation, we backpropagate the update term along the prior state chain and asymptotically compensate for residual distortion before the next iteration. Second, considering the weak correlation between the initial error and motion distortion in the current frame, we propose a convergence criteria based on pointcloud constraints to control backpropagation. In summary, the approach of guiding asymptotic compensation based on convergence criteria is designed to promote the consistent convergence of pointcloud registration, ensuring robustness while further improving the odometry accuracy.
We conducted a series of experiments on public and self-collected datasets, covering various scenarios and motion conditions across multiple scales. The experiments show that, compared to other state-of-the-art frameworks, our AC-LIO framework effectively promotes consistent convergence in state estimation and further enhances the accuracy of long-term, large-scale localization and mapping.

### v2

@@
 Abstract
-Existing LiDAR-Inertial Odometry (LIO) frameworks typically utilize prior state trajectories derived from IMU integration to compensate for the motion distortion within LiDAR frames, and demonstrate outstanding accuracy and stability in regular low-speed and smooth scenes.
-However, in high-speed or intense motion scenarios, the residual distortion may increase due to the limitation of IMU’s accuracy and frequency, which will degrade the consistency between the LiDAR frame with its represented geometric environment, leading pointcloud registration to fall into local optima and consequently increasing the drift in long-time and large-scale localization.
-To address the issue, we propose a novel asymptotically and consistently converging LIO framework called AC-LIO.
-First, during the iterative state estimation, we backwards propagate the update term based on the prior state chain, and asymptotically compensate the residual distortion before next iteration.
-Second, considering the weak correlation between the initial error and motion distortion of current frame, we propose a convergence criteria based on pointcloud constraints to control the back propagation.
-The approach of guiding the asymptotic distortion compensation based on convergence criteria can promote the consistent convergence of pointcloud registration and increase the accuracy and robustness of LIO.
-Experiments show that our AC-LIO framework, compared to other state-of-the-art frameworks, effectively promotes consistent convergence in state estimation and further improves the accuracy of long-time and large-scale localization and mapping.
+Existing LiDAR-Inertial Odometry (LIO) methods typically utilize the prior state trajectory derived from the IMU integration to compensate for the motion distortion within LiDAR frames.
+However, discrepancies between the prior and actual trajectory can lead to residual distortions that compromise the consistency of the LiDAR frame with its corresponding geometric environment.
+This imbalance may result in pointcloud registration becoming trapped in local optima, thereby exacerbating drift during long-term and large-scale localization.
+To address the issue, we propose a novel asymptotically and consistently converging LIO framework dubbed AC-LIO.
+Our key idea is to back propagate current update term based on the prior state chain, and asymptotically compensate for the residual distortion during iteration.
+Moreover, considering the weak correlation between previous error and current distortion, we establish convergence criteria based on the pointcloud constraints to regulate the backpropagation.
+This method of guiding asymptotic distortion compensation using convergence criteria subtly enhances the consistent convergence of pointcloud registration, futher improving the accuracy and robustness of LIO system.
+Extensive experiments demonstrate that our AC-LIO framework significantly promotes consistent convergence in state estimation compared to prior arts, with about 30.4% reduction in average RMSE over the second best result, leading to marked improvements in the accuracy of long-term and large-scale localization and mapping.
 I INTRODUCTION
-Figure 1: The basic principle of AC-LIO framework.
-Conventional LIO typically consider initial distortion compensation via IMU integration, yet the residual distortion may prevent the registration from further convergence.
-In contrast, our AC-LIO employs an asymptotic backpropagation strategy to promote consistent convergence in state estimation.
-With the development of Simultaneous Localization and Mapping (SLAM) technology, the perception of robots on the external environment and their own motion is becoming more and more precise [1, 2].
-As a key branch, LiDAR-Inertial Odometry (LIO) has gained widespread application in autonomous driving, unmanned logistics and other fields, for its high robustness and accuracy [3, 4].
-Compared to Visual-Inertial Odometry (VIO), the LIO system can directly obtain depth observations of surroundings [5], without the need to estimate the depth of feature points [6, 7], and exhibits better adaptability to different lighting conditions [8, 9].
-However, as LIO system continues to evolve, there still remain several fundamental issues to be addressed [10, 11], one of which is the conflict between continuous motion and rigid pointcloud registration [12].
-Since the acquisition of LiDAR stream is often accompanied with the ego-motion of vehicle, the laser points with different timestamps actually lie in different coordinate frames.
-Existing frameworks typically only consider using IMU integration for the initial compensation of motion distortion [13, 14, 15], without further optimization of the residual distortion in state estimation.
-But in high-speed or violent motion scenarios, the residual motion distortion tends to increase due to the limitations of IMU’s accuracy and frequency, which would suppress further convergence of pointcloud registration and decrease odometry accuracy in long-term, large-scale scenarios [16].
-To address this, we propose AC-LIO, a novel asymptotically and consistently converging LIO framework.
-First, during the iterative state estimation, we backwards propagate the current update term based on the prior state chain, and asymptotically compensate the residual distortion before next iteration.
-This approach could enhance the consistency between the current LiDAR frame and the represented geometric environment, and avoid the pointcloud registration from falling into local optimum due to the residual distortion, thereby promoting the consistent convergence of state estimation and improving the accuracy of long-time and large-scene localization.
-Compared to other strategies such as continuous-time estimation or frame segmentation, our method can be seamlessly embedded into the iterative Kalman Filter framework [17], offering great flexibility and scalability.
-Second, considering the weak correlation between the initial error and motion distortion of current frame, we also propose a convergence criteria based on the point-plane residuals to control the back propagation.
-This strategy ensures, We use this criteria to evaluate the convergence of the pointcloud registration, ensuring the backpropagation and asymptotic compensation only happens under the condition of sufficient convergence in previous LiDAR frame, and thus guaranteeing the robustness of LIO system.
+With the advancements in Simultaneous Localization and Mapping (SLAM) technology, robots are increasingly achieving precise perception of their external environments and their own motion[1, 2, 3].
+As a vital subset of SLAM, LiDAR-Inertial Odometry (LIO) has gained widespread application in autonomous driving [4], unmanned logistics [5], humanoid robot [6] and other fields, for its high robustness and accuracy [7, 8].
+Compared to Visual-Inertial Odometry, the LIO system can directly obtain depth observations of surroundings [9], without the need to estimate the depth of feature points [10, 11], and exhibits better adaptability to different lighting conditions [12, 13].
+Figure 1: (a) A principal comparison of our AC-LIO with traditional approaches.
+Conventional LIO typically considers an initial distortion compensation via IMU, yet the residual distortion may prevent the registration from further convergence.
+In contrast, our AC-LIO employs an asymptotic backpropagation to promote consistent convergence in state estimation.
+(b) Large scale localization and mapping result of BotanicGarden dataset.
+We achieve better accuracy performance than other benchmark frameworks on its sequence 1008_01 with VLP16 LiDAR.
+However, as LIO system continues to evolve, there still remain several fundamental issues to be addressed [14, 15, 16].
+One of them is the conflict between continuous motion and rigid pointcloud registration [17].
+Since the acquisition of LiDAR frames is often accompanied with the ego-motion of the vehicle, the LiDAR points with different timestamps actually lie in different coordinate systems.
+As shown in Fig.
+1, traditional methods [18, 19, 20] typically only consider using IMU integration for the initial compensation of motion distortion, without further optimization of the residual distortion in the state estimation.
+This residual distortion will compromise the consistency between the LiDAR frame with its corresponding geometric environment, and it grows with increasing motion non-linearity.
+Therefore, in conventional LIO systems, such residual error may suppress further convergence of pointcloud registration during the iterative update, ultimately impacting the odometry accuracy and causing severe drift.
+In this letter, we propose AC-LIO, a novel asymptotically and consistently converging LIO framework.
+Our AC-LIO enjoys two key technical contributions.
+First, during the iterative state estimation, we introduce an asymptotic distortion compensation (Sec.
+III-C) method that backwards propagate the current update term based on the prior state chain, and asymptotically compensate the residual distortion before next iteration .
+This approach could enhance the consistency between the current LiDAR frame and the represented geometric environment, and avoid the pointcloud registration from falling into local optimum due to the residual distortion, thereby promoting the consistent convergence of pointcloud registration and improving the accuracy of long-term and large-scene localization.
+Second, considering the weak correlation between initial error and motion distortion of current frame, we also propose the convergence criteria (Sec.
+III-D) based on the point-to-plane residuals to control the back propagation.
+The criteria is used to assess the convergence of the pointcloud registration and ensure that the backpropagation and asymptotic compensation only happens under the condition of sufficient convergence in previous LiDAR frame, hence guaranteeing the robustness of LIO system.
+Compared to other strategies such as continuous-time estimation or frame segmentation, our method leverages this criteria to construct a feedback mechanism that promotes state convergence without additional estimation, achieving a balance between localization accuracy and computational efficiency (Sec.
+IV-C).
+It can be seamlessly embedded into the iterative Kalman Filter framework [21], offering great flexibility and scalability.
 In summary, our main contributions are as follows:
-  • We implement a novel asymptotically and consistently converging LIO framework.
-During the iterative update, we back propagate the current update term based on the prior state chain, and asymptotically compensate the residual distortion before next iteration, promoting the consistent convergence of registration and improving the accuracy in long-time and large-scale scenarios.
-  • We introduce a convergence criteria based on point-plane residuals to control the back propagation.
-We utilize the criteria to evaluate the convergence of registration, ensuring the back propagation to be performed only upon sufficient convergence of the previous registration, thus further enhancing the system robustness.
-  • Our method can be easily and seamlessly integrated into traditional iterative Kalman Filter framework.
-Compared to other optimization strategies for motion distortion, our framework introduce no extra estimates or states, making the asymptotic compensation less computationally expensive but highly flexible and scalable.
+  • We propose a novel framework AC-LIO with asymptotic distortion compensation that showing plausible consistent convergence in long-term and large-scale scenarios.
+  • We introduce the convergence criteria based on point-to-plane residuals to evaluate the convergence and guide back propagation.
+Back propagation is only performed upon sufficient convergence of previous registration, thus further enhancing the system robustness.
+  • Compared to other optimization strategies for motion distortion, our framework introduce no extra estimates or states, making the asymptotic compensation less computationally expensive but highly flexible.
   • We conduct extensive experiments on AC-LIO in a series of real-world datasets with various scenarios.
-The results show that our proposed AC-LIO exhibits higher accuracy in long-term localization compared to the state-of-the-art open-source methods.
+The results show that our method achieves higher accuracy compared to current state-of-the-art frameworks.
 II RELATED WORK
-Since the motion distortion can degrade the consistency between the pointcloud and the geometric environment, and thus suppress the sufficient convergence of registration.
-To mitigate the adverse effect of motion distortion, existing LiDAR(-inertial) odometry adopt various strategies to appropriately compensate for this distortion.
-For LiDAR-only odometry, most studies utilize the constant velocity model to compensate for motion aberrations.
-Studies such as LOAM [18] and Lego-LOAM [19] have achieved significant effects with the assumption of constant velocity motion during the corresponding time of LiDAR frames.
+LiDAR-Only Odometry.
+Studies such as LOAM [22] and F-LOAM [23] has achieved significant performance with the assumption of constant velocity motion during the corresponding time of LiDAR frames.
 However, this assumption only works effectively when LiDAR frame interval is enough short and the motion is slow and smooth.
 In case of drastic velocity variation, the constant velocity model cannot accurately characterize the current motion state, and this will result in increased drift and even odometry divergence.
-As for LiDAR-inertial odometry, the IMU observation can provide relatively reliable prior state information.
-This provides a new scheme for compensating motion distortion via IMU integration.
-FAST-LIO [20] employs the IMU state backpropagation to correct the distorted pointcloud, and then conducts state estimation through Error State Kalman Filter.
-LIO-SAM [21] utilizes the IMU pre-integration to de-skew pointcloud, and combines multiple constraints including LiDAR, IMU, and GPS through the factor graph.
+CT-ICP [24] optimize the start and end states of LiDAR frame, and construct a continuous time trajectory by interpolation.
+But the real movement may not follow the interpolated trajectory, as well as the logical constraint across adjacent frames is hard to determine.
+I2EKF-LO [25] utilizes extra prediction iteration to estimate motion distortion, yet the update of the process noise still depends on the hyperparameters.
+Traj-LO [26] leverages the high frequency of LiDAR points and employs more linear segments within frame.
+This approach allows more accurate trajectory characterization and suppression of motion aberrations during high speed motion, but also introduces more computational cost.
+Discrete State LiDAR-Inertial Odometry.
+For LiDAR-inertial odometry, the IMU observation can provide relatively reliable prior state information.
+FAST-LIO2 [27] employs the IMU state backpropagation to correct the distorted pointcloud, and then conducts state estimation through Error State Kalman Filter.
+LIO-SAM [28] utilizes the IMU pre-integration to de-skew pointcloud, and fuses multiple constraints including LiDAR, IMU, and GPS through the factor graph.
+In the subsequent LIO frameworks, SR-LIO [29] adopts a LiDAR frame reconstruction strategy, while AS-LIO [30] designs an adaptive sliding window guided by spatial overlap degree, both of which can boost the system update frequency and lower the impact of motion distortion.
+Semi-Elastic-LIO [31] replaces the interpolation model with IMU motion constraints on the basis of CT-ICP framework, thereby improving the stability of the inter-frame logical constraints.
 Although these methods can mitigate the impact of pointcloud distortion to some extent, there are still some limitations without considering the further optimization of residual motion distortion.
+Point-LIO [32] adopts a point-by-point update strategy to avoid the motion distortion issue and realize extremely high-frequency odometry.
+Compared to frame based registration, the point-by-point strategy is more likely to affected by dynamic objects and degenerate in adverse scenarios.
+Continuous State LiDAR-Inertial Odometry.
 To characterize trajectorie more precisely in state estimation, many frameworks come up with new optimization solutions.
-As pure LiDAR framework, I2EKF-LO [22] utilizes additional iteration to specifically estimate motion distortion, this approach can suppress distortion but the extra iteration may sacrifice computational efficiency.
-CLINS [23] introduces the B-spline model [24] for continuous time state estimation.
-However, the unconstrained trajectory would not always conform to the B-spline model due to time resolution limitations in high-speed, dynamic motion.
+CLINS [33] introduces the B-spline [34] for continuous time state estimation, while SLICT [35] employs continuous-time factors to represent LiDAR constraints within a sliding window.
+These frameworks further optimize motion distortion with continuous-time model.
+However, the unconstrained trajectory would not always conform to the predefined model due to time resolution limitations in high-speed, dynamic motion.
 Moreover, the continuous-time optimization strategy may introduce a higher computational cost.
-Point-LIO [25] adopts a point-by-point update strategy to avoid the motion distortion issue and realize extremely high-frequency odometry.
-Compared to frame based registration, the point-by-point strategy is more likely to degenerate in adverse scenarios.
-FR-LIO [26] divides the LiDAR frame into several sub-frames based on IMU observation.
-This approach can increase the system update frequency, thus reducing the impact of motion distortion.
-Besides, FR-LIO strengthens the constraints between subframes via RTS smoother [27], but still faces the risk of subframe degradation.
+DLIO [36] uses analytical equations to construct a continuous trajectory for precise motion correction, after coarse deskewing through IMU integration.
+But the method remains limited by IMU’s inherent accuracy and neglects further undistortion during the registration process.
 In contrast, our proposed AC-LIO framework utilizes the convergence criteria to guide the backpropagation process, to achieve the asymptotic compensation of residual motion distortion.
@@
 III METHODOLOGY
+TABLE I: Important Notations
+Symbols Definition
+[…]
+$I,L,G$ The IMU, LiDAR, and global coordinate frame.
+$\mathbf{x}^{gt}$ The ground truth of state $\mathbf{x}$.
+$\mathbf{x}_{i}$ The vector of state $\mathbf{x}$ at timestamp $i$.
+$\hat{\mathbf{x}}$ The priori estimation of state $\mathbf{x}$.
+$\hat{\mathbf{x}}^{j},\delta\hat{\mathbf{x}}^{j}$ The $j$-th update of state $\mathbf{x}$ and error state $\delta\mathbf{x}$.
+$\bar{\mathbf{x}},\check{\mathbf{x}}$ The updated state $\mathbf{x}$ with or without back propagation.
+$\mathbf{F_{a}}^{i|i+n}$ The propagation matrix from state $i$ to $i+n$.
+$\mathbf{G}_{i}^{i+n}$ The backward gain matrix of state $i$ with respect to $i+n$.
+$\mathbf{z},z$ The vector and its scalar value of point-to-plane residual.
+$\phi$ The incident angle of single laser beam on the fitted plane.
+$\sigma_{t}$ The expectation of average point-to-plane residual.
 III-A System Overview
-Figure 2: The system overview of our AC-LIO framework.
-During iterative ESKF, asymptotic compensation of pointcloud distortion will be conducted based on the convergence criteria, to promote consistent convergence of state estimation.
+Figure 2: (a) The system overview of our AC-LIO framework.
+During iterative ESKF, asymptotic compensation of pointcloud distortion is conducted based on the convergence criteria, to promote consistent convergence of state estimation.
+(b) The on-manifold propagation.
+The blue shows the forward propagation of IMU integration.
+The green indicates the backward recursion of the update term according to the propagation matrix and noise relation.
 The system overview of our AC-LIO framework is shown in Fig.
-2.
-IMU observations are fed into the forward propagation module for integration to obtain prior trajectories.
-Then, the LiDAR data utilize the prior state chain for initial compensation, next passed into the Error State Kalman Filter (ESKF) for state estimation [28], where the average point-to-plane residual (APR) is calculated after each update.
+2, and the important notations is shown in Table I.
+First, the IMU observations are fed into the forward propagation module to obtain prior trajectories.
+Then the LiDAR data utilize the prior state chain for initial compensation and then sent to the Error State Kalman Filter (ESKF) for state estimation [37], where the average point-to-plane residual (APR) is calculated after each update.
 During the iteration, if the APR of previous LiDAR frame is lower than the convergence criteria and the current APR exceeds it, then the backpropagation of update residual will be executed.
-After backpropagation, the LiDAR frame needs to be re-compensated to promote the consistent convergence of state estimation.
+The LiDAR frame will be re-compensated to promote the convergence of registration after each backpropagation.
 Otherwise, if the conditions are not met, the backpropagation will be skipped and iteration continues.
@@
 Based on the kinematic model and IMU measurement model, the state $\mathbf{x}$ , input $\mathbf{i}$, and noise $\mathbf{n}$ can be defined as follows:
-$\begin{split}\mathbf{x}&\doteq\begin{bmatrix}\mathbf{R}_{I}^{G}&\mathbf{p}_{I}^{G}&\mathbf{v}_{I}&\mathbf{b}_{\bm{\omega}}&\mathbf{b}_{\mathbf{a}}&\mathbf{g}^{G}\end{bmatrix}^{T}\\ \mathbf{i}&\doteq\begin{bmatrix}{\bm{\omega}}_{m}&{\mathbf{a}}_{m}\end{bmatrix}^{T}\\ \mathbf{n}&\doteq\begin{bmatrix}\mathbf{n}_{\bm{\omega}}&\mathbf{n}_{\mathbf{a}}&\mathbf{n}_{\mathbf{b}\bm{\omega}}&\mathbf{n}_{\mathbf{b}\mathbf{a}}\end{bmatrix}^{T}\end{split}$ (1)
-Where $\mathit{G}$ and $\mathit{I}$ denote the global and the IMU frames, $\mathbf{R}_{I}^{G}$, $\mathbf{p}_{I}^{G}$ and $\mathbf{v}_{I}^{G}$ are IMU’s orientation, position and linear velocity in $\mathit{G}$, and $\mathbf{g}^{G}$ is the local gravity.
-The $\mathbf{n}_{\bm{\omega}}$, $\mathbf{n}_{\mathbf{a}}$ are noises of angular velocity ${\bm{\omega}}_{m}$ and linear acceleration ${\mathbf{a}}_{m}$, the $\mathbf{b}_{\bm{\omega}}$, $\mathbf{b}_{\mathbf{a}}$ are IMU bias with $\mathbf{n}_{\mathbf{b}\bm{\omega}}$ and $\mathbf{n}_{\mathbf{b}\mathbf{a}}$ as their noises in Gaussian-Markov model.
+$\begin{split}\mathbf{x}&\doteq\begin{bmatrix}\mathbf{R}_{I}^{G}&\mathbf{p}_{I}^{G}&\mathbf{v}_{I}&\mathbf{b}_{\bm{\omega}}&\mathbf{b}_{\mathbf{a}}&\mathbf{g}^{G}\end{bmatrix}^{T}\\ \mathbf{i}&\doteq\begin{bmatrix}{\bm{\omega}}_{m}&{\mathbf{a}}_{m}\end{bmatrix}^{T},~{}~{}\mathbf{n}\doteq\begin{bmatrix}\mathbf{n}_{\bm{\omega}}&\mathbf{n}_{\mathbf{a}}&\mathbf{n}_{\mathbf{b}\bm{\omega}}&\mathbf{n}_{\mathbf{b}\mathbf{a}}\end{bmatrix}^{T}\end{split}$ (1)
 […]
 $\begin{gathered}\quad\quad\delta\mathbf{x}_{i}\doteq\begin{bmatrix}\delta{\bm{\theta}}&\delta\mathbf{p}_{I}^{G}&\delta\mathbf{v}_{I}&\delta\mathbf{b}_{\bm{\omega}}&\delta\mathbf{b}_{\mathbf{a}}&\delta\mathbf{g}^{G}\end{bmatrix}^{T}\\ \quad\delta\hat{\mathbf{x}}_{i+1}=\mathbf{F_{a}}\delta\check{\mathbf{x}}_{i}+\mathbf{F_{n}}\mathbf{n}_{i}\\ \quad\hat{\mathbf{P}}_{i+1}=\mathbf{F_{a}}\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}+\mathbf{F_{n}}\mathbf{Q}\mathbf{F_{n}}^{T}\end{gathered}$ (2)
-Denote $\mathbf{T}_{L}^{I}$ is the extrinsic between LiDAR and IMU, then the pointcloud can be transformed to frame $\mathit{I}$ via the prior pose chain.
+Assume the extrinsic $\mathbf{T}_{L}^{I}$ between LiDAR and IMU is already known.
 For a point $\mathbf{p}$ in LiDAR frame, its neighboring cluster can be searched by KNN in pointcloud map.
-A plane is then fitted to this cluster and we obtain the normal vector $\mathbf{u}$.
-Note one point of the cluster as $\mathbf{q}$ and $(\cdot)_{\wedge}$ as the upper triangular matrix of vector, then the point-plane residual $\mathbf{z}$ and measurement matrix $\mathbf{H}$ are as follows:
+If a plane $S$ can be fitted to this cluster, then we will obtain its normal vector $\mathbf{u}$.
+Note one point of the cluster as $\mathbf{q}$ and $(\cdot)_{\wedge}$ is the upper triangular matrix of vector, then the point-to-plane residual $\mathbf{z}$ and measurement matrix $\mathbf{H}$ are as follows:
 $\begin{gathered}\mathbf{z}=\mathbf{u}(\hat{\mathbf{p}}^{G}-\mathbf{q}^{G})^{T}\\ \mathbf{H}=\begin{bmatrix}-\mathbf{u}\hat{\mathbf{R}}_{I}^{G}(\mathbf{T}_{L}^{I}\mathbf{p}^{L})_{\wedge}&\mathbf{u}&\mathbf{0}_{3\times 12}\end{bmatrix}\end{gathered}$ (3)
@@
 III-C Asymptotic Distortion Compensation
-In smooth motion scenarios, an initial compensation of motion distortion via IMU integration can basically satisfy the accuracy demand.
-However, when encountering violent motion scenarios, the residual motion distortion may increase due to the limitation of IMU’s frequency and accuracy.
 Compared to the noise of LiDAR observations, motion distortion may introduce bias in pointcloud registration, making it difficult to utilize the statistical properties of pointcloud to mitigate the adverse effects on state estimation.
-With reference to the fixed interval smooth strategy, we selectively back propagate the update term $\delta\mathbf{x}$ based on the prior state chain during iteration.
-The noise between adjacent states can be expressed as follows:
-$\begin{gathered}\delta\check{\mathbf{x}}_{i}\sim\mathcal{N}(\mathbf{0}_{18\times 1},\check{\mathbf{P}}_{i})\\ \delta\hat{\mathbf{x}}_{i+1}\sim\mathcal{N}(\mathbf{0}_{18\times 1},\mathbf{F_{a}}\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}+\mathbf{F_{n}}\mathbf{Q}\mathbf{F_{n}}^{T})\end{gathered}$ (6)
-Figure 3: The on-manifold propagation.
-The noise of current state involves the process noise and the propagation of previous state, the update term can be recursive backward according to this relationship.
 During the on-manifold propagation of state $\mathbf{x}$, the noise of current error state $\delta\hat{\mathbf{x}}_{i+1}$ consists of the propagation of previous state $\delta\hat{\mathbf{x}}_{i}$ and the accumulated process noise.
-According to RTS smooth strategy [27], the contribution of previous state in current update term just corresponds to the proportion of the propagation covariance $\mathbf{F_{a}}\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}$ within current covariance $\hat{\mathbf{P}}_{i+1}$.
-Consequently, the gain matrix $\mathbf{G}$ and the backward recursion of $\delta\mathbf{x}$ can be denoted as follows:
+With reference to the Rauch-Tung-Striebel fixed interval smooth strategy [39], we can back propagate the update term $\delta\mathbf{x}$ based on the prior state chain.
+The contribution of previous state in current update term just corresponds to the proportion of propagation covariance $\mathbf{F_{a}}\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}$ within current covariance $\hat{\mathbf{P}}_{i+1}$.
+Then, the gain matrix $\mathbf{G}$ and the backward recursion of $\delta\mathbf{x}$ can be denoted as follows:
 $\begin{gathered}\mathbf{G}=\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}\hat{\mathbf{P}}_{i+1}^{-1}\\ \delta\hat{\mathbf{x}}_{i}^{j}=\mathbf{G}\cdot\delta\hat{\mathbf{x}}_{i+1}^{j}\end{gathered}$ (6)
-Furthermore, since there may exist numerous IMU states to maintain within the timeperiod of LiDAR frame, a sequential backward recursion can characterize the smoothed trajectory precisely.
-However, from an overall perspective, this treatment will increase much computational cost in iterative ESKF, but cannot bring additional accuracy improvement in subsequent asymptotic distortion compensation.
-Compared to state-by-state backpropagation, selectively updating several anchors in the prior state chain can further reduce the computational cost with little accuracy loss.
-To this end, we extend the backwards recursion to any states within the LiDAR frame, thereby reducing the computational complexity of backpropagation and distortion compensation.
+Furthermore, this sequential backpropagation method can be extend to any states within the LiDAR frame.
+It allows us to selectively update several anchor in the prior state chain, thus reducing the computational complexity of backpropagation and distortion compensation.
+As for any states $\mathbf{x_{i}}$ and $\mathbf{x_{i+n}}$ in the prior state chain, the propagation matrix and backward gain matrix can be noted as $\mathbf{F_{a}}^{i|i+n}$ and $\mathbf{G}_{i}^{i+n}$, then the backpropagation relationship should be:
 $\begin{gathered}\mathbf{F_{a}}^{i|i+n}=\prod_{k=i}^{i+n-1}\mathbf{F_{a}}^{k|k+1}\\ \mathbf{G}_{i}^{i+n}=\check{\mathbf{P}}_{i}(\mathbf{F_{a}}^{i|i+n})^{T}(\hat{\mathbf{P}}_{i+n})^{-1}\\ \check{\mathbf{x}}_{i}^{j}=\check{\mathbf{x}}_{i}\boxplus(\mathbf{G}_{i}^{i+n}\cdot\delta\hat{\mathbf{x}}_{i+n}^{j})\end{gathered}$ (7)
@@
 III-D Convergence Criteria
-In order to exploit the statistical properties of the pointcloud, during the iterative state estimation, we select the mean value of the point-to-plane residual (abbreviated as APR above), as the criteria for evaluating the convergence of pointcloud registration.
-For simplification, assuming the pointcloud registration fully converges to the global optimum, considering the ranging accuracy of LiDAR and the construction of point-to-plane residual, we can roughly model the expectation of the APR, to provide a threshold reference for the convergence criteria.
-Denote the fitted plane in pointcloud map as $S$ with its normal vector as $\mathbf{u}$, and the point-plane residual as $\mathbf{z}$.
-Then, the incident angle $\phi$ of point $\mathbf{p}$ with respect to plane $S$ can be expressed as:
-$\begin{gathered}\phi=\arccos\dfrac{|\mathbf{u}\cdot\mathbf{p}|}{|\mathbf{u}|\cdot|\mathbf{p}|}\end{gathered}$ (11)
-When the laser beam is perpendicular to plane $S$, then the incident angle $\phi$ will be zero.
-Assuming that the ranging error $\mathbf{n_{p}}$ of laser point follows Gaussian distribution with zero mean and $\sigma$ variance, then the residual $\mathbf{z}_{\perp}$ will follow chi-squared distribution and can be expressed as follows:
-$\begin{gathered}\mathbf{p}^{L}=\mathbf{p}_{gt}^{L}+\mathbf{n_{p}},\;\mathbf{n_{p}}\sim\mathcal{N}(0,\sigma)\\ \mathbf{z}_{\perp}=|\mathbf{p}_{gt}^{L}-\mathbf{p}^{L}|\\ \frac{\mathbf{z}_{\perp}^{2}}{\sigma^{2}}\sim\chi^{2}(1),\;\mathbb{E}(\mathbf{z}_{\perp})=\sigma\end{gathered}$ (12)
-As the surrounding environment varies during ego-motion, the distribution of normal vector $\mathbf{n}$ of plane $S$ will also change.
-Eventually, the variation of the geometrical scene will be reflected in the distribution of the incident angle $\phi$.
-Assuming the incident angle is uniformly distributed over the domain $[0,\pi/2)$ , then the expected mean of the point-plane residual $\mathbf{z}$ can be expressed as:
-$\begin{gathered}\mathbb{E}(\mathbf{z})=\frac{2}{\pi}\int_{0}^{\frac{\pi}{2}}\mathbb{E}(\mathbf{z}_{\perp})\cos{\phi}\;d\phi=\frac{2}{\pi}\sigma\end{gathered}$ (13)
+Figure 3: The relationship between original ranging variance of LiDAR point and point-to-plane residual variance, where the residual $z$ is a scalar.
+To exploit the statistical properties of pointcloud, during the iterative state estimation, we select the mean value of the point-to-plane residual (abbreviated as APR above), as the criteria for evaluating the convergence of pointcloud registration.
+For simplification, assuming the pointcloud registration fully converges to the global optimum and only considering the ranging accuracy of LiDAR, we can roughly model the expectation of the APR and provide a threshold reference for the convergence criteria.
+Denote the incident angle of point $\mathbf{p}$ to the fitted plane $S$ as $\phi$, it can be expressed as $\cos^{-1}\tfrac{|\mathbf{u}\cdot\hat{\mathbf{p}}^{G}|}{|\mathbf{u}|\cdot|\hat{\mathbf{p}}^{G}|}$ according to Eq.
+3.
+Note that $\mathbf{z}$ is the point-to-plane residual vector and $z$ is its scalar value.
+Assuming the ranging error of LiDAR point follows Gaussian distribution with zero mean and $\sigma$ variance.
+When the laser beam is perpendicular to the fitted plane, the residual $z_{\perp}$ follows chi-squared distribution and can be expressed as follows:
+$\begin{gathered}{z}_{\perp}=|\mathbf{p}^{G}-\hat{\mathbf{p}}^{G}|\sim\mathcal{N}(0,\sigma)\\ \frac{{z}_{\perp}^{2}}{\sigma^{2}}\sim\chi^{2}(1),\;\mathbb{E}({z}_{\perp})=\sigma\end{gathered}$ (10)
+As the surroundings vary during system’s ego-motion, the distribution of normal vector $\mathbf{u}$ to plane $S$ also changes.
+The variation of the geometrical scene would be reflected in the distribution of the incident angle $\phi$.
+From Fig.
+3, the residual variance $\sigma_{z}$ should be $\sigma\cos\phi$.
+Assuming the incident angle is uniformly distributed over the domain $[0,\pi/2)$, then the expected mean of point-to-plane residual can be expressed as:
+$\begin{gathered}\mathbb{E}({z})=\frac{2}{\pi}\int_{0}^{\frac{\pi}{2}}\mathbb{E}({z}_{\perp})\cos{\phi}\;d\phi=\frac{2}{\pi}\sigma\end{gathered}$ (11)
 Therefore, assuming the pointcloud registration converges to the global optimum, and only considering the LiDAR ranging error, the expectation of APR should be ${2\sigma}/{\pi}$.
-If the mean value of current point-plane residual is lower than this criteria, then the last registration can be considered as sufficiently converged.
-According to the threshold of convergence criteria, we can easily determine the convergence degree of pointcloud registration, and control the backpropagation and the asymptotic distortion compensation.
-Noting that $\sigma_{t}=\mathbb{E}(\mathbf{z})$ and the selected threshold of convergence criteria as $\eta$, due to the factors such as dynamic objects and cumulative errors in pointcloud map, the threshold $\eta$ can be appropriately extended to the interval of $[\sigma_{t},2\sigma_{t}]$ in practice.
-In particular, we perform two conditional judgments in iterative state estimation: First, evaluate the previous LiDAR frame, if the APR is less than the threshold, it is considered to be sufficiently converged and we can continue.
-Then, check the convergence of current registration, if the APR is still less than the threshold, there is no need to back-propagate for small residual distortion; otherwise, perform the backpropagation and asymptotic compensation, in order to promote the consistent convergence in state estimation.
-The approach, on one hand, can ensure the backpropagation and asymptotic compensation only performed as the previous LiDAR frame converged well, thus enhancing the odometry robustness in long-time, large-scale scenario; on the other hand, can avoid lots of additional operations in unreliable or unnecessary conditions, further saving the system computational expense.
+If the mean value of current point-to-plane residual is lower than this criteria, the last registration can be regarded as sufficiently converged.
+Based on this threshold of convergence criteria, we can easily determine the convergence degree of pointcloud registration, and control the backpropagation and asymptotic distortion compensation.
+Noting that $\sigma_{t}=\mathbb{E}({z})$ and the selected threshold of convergence criteria as $\eta$, due to the factors such as dynamic objects and cumulative errors in pointcloud map, the threshold $\eta$ can be appropriately extended to the interval of $[\sigma_{t},2\sigma_{t}]$ in practice.
+In detail, we perform two conditional judgments in iterative state estimation: First, evaluate the previous LiDAR frame, if its APR is less than the threshold $\eta$, it can be considered to be sufficiently converged and we will continue.
+Then, check the convergence of current registration, if the APR is still less than $\eta$, there is no need to backpropagate for small residual distortion; otherwise, we will perform the backpropagation and compensation during the iteration.
+On one hand, this approach can ensure the backpropagation and asymptotic compensation only performed as the previous LiDAR frame converged well, thus enhancing the odometry robustness in long-term and large-scale scenario.
+On the other hand, it can avoid lots of additional operations in unreliable or unnecessary conditions, to improve the system computational efficiency.
 The experiment results of our framework are shown in the next section.
 IV EXPERIMENT RESULTS
+TABLE II: Odometry Accuracy Evaluation - RMSE of ATE & End to End Errors($m$)
+Dataset LiDAR Type Sequence Ours FAST-LIO2 DLIO SLICT Point-LIO
+WHU1 Helmet Livox AVIA Residence_5.2 (452.9m) 0.415 0.505 4.660 1.550 $\times$
+Park_3.2 (628.9m) 0.582 0.618 7.725 0.602 $\times$
+Subway_3.3 (827.0m) 1.670 3.752 3.688 1.967 $\times$
+Mall_6.3 (1006.5m) 0.216 0.208 3.238 0.274 15.409
+Botanic1 Garden Livox AVIA 1005_05 (566.8m) 0.773 1.053 0.640 7.931 0.917
+1006_03 (686.2m) 1.796 2.075 3.017 6.527 2.471
+1008_01 (855.9m) 1.152 1.577 14.829 4.445 1.590
+VLP16 1005_05 (566.8m) 0.175 0.244 0.565 37.425 0.750
+1006_03 (686.2m) 0.377 0.388 0.393 $\times$ 0.445
+1008_01 (855.9m) 0.339 0.363 0.694 $\times$ 0.457
+Handheld2 Livox AVIA Indoor_1 (~100m) 0.893 1.133 $\times$ 1.960 5.401
+Indoor_2 (~100m) 0.227 0.204 3.354 0.249 0.164
+Indoor_3 (~100m) 1.596 3.001 $\times$ 2.476 6.844
+Outdoor_1 (~300m) 0.138 1.901 0.486 1.835 0.728
+Outdoor_2 (~400m) 0.984 2.194 7.879 1.803 1.261
+Outdoor_3 (~500m) 2.666 4.702 1.445 4.127 4.607
+Average(RMSE) 0.750 1.078 3.945 7.590 3.148
+  • 1  means the dataset is evaluated by RMSE of ATE, and  2 means it is evaluated by End to End errors.
+  • $\times$  denotes the system severely diverged midway, and $\underline{\hskip 7.11317pt}$  denotes the second best result compared to bolded result.
 In this section, we qualitatively and quantitatively evaluate AC-LIO framework through a series of experiments with various public and self-collected datasets.
-To examine the accuracy improvement of our approach in long-time and large-scale scenarios, and the effect of convergence criteria on system robustness, we conduct extensive experiments on the odometry accuracy and the validity of convergence criteria.
-IV-A Odometry Accuracy Evaluation
-In the odometry accuracy experiments, we compare AC-LIO framework with current state-of-the-art LIO systems including FAST-LIO2 and Point-LIO.
-All algorithms are implemented in C++ and employed on an i9 CPU desktop computer with the ROS operating system in Ubuntu.
-These methods only utilize CPU for computation without the assistance of GPU parallel computing.
-Besides, in the odometry accuracy experiments, we maintain the threshold of convergence criteria in the extended interval, i.e., $\eta\in[\sigma_{t},2\sigma_{t}]$.
-A brief introduction of the public and self-collected handheld datasets is provided below:
+To examine the performance of our approach in long-term and large-scale scenarios, we conduct extensive experiments on the odometry accuracy, the time efficiency of system, and the validity of convergence criteria.
 Figure 4: Our handheld sensor suite used in self-collected dataset, including the solid-state LiDAR Livox Avia with the built-in BMI088 IMU, and three other cameras.
-TABLE I: Odometry Accuracy Evaluation - RMSE of ATE ($m$)
-AC-LIO(Ours) FAST-LIO2 Point-LIO
-Residence_5.2
-0.415 0.505 1
-Park_3.2
-0.582 0.618
-Subway_3.3
-1.670 3.752
-Mall_6.3
-0.216 0.208 15.409
-Garden_1005
-0.773 1.053 0.917
-Garden_1006
-1.796 2.075 2.471
-Garden_1008
-1.152 1.577 1.590
-  • 1
-    –  means that the system severely diverged midway.
-TABLE II: Odometry Accuracy Evaluation - End to end errors ($m$)
-AC-LIO(Ours) FAST-LIO2 Point-LIO
-Indoor_1
-(~100m)
-0.8929 1.1330 5.4006
-Indoor_2
-(~100m)
-0.2267 0.2043 0.1636
-Indoor_3
-(~100m)
-1.5959 3.0011 6.8438
-Outdoor_1
-(~300m)
-0.1378 1.9007 0.7280
-Outdoor_2
-(~400m)
-0.9840 2.1944 1.2607
-Outdoor_3
-(~500m)
-2.6659 4.7023 4.6066
-  • WHU-Helmet [30]: Collected on compact helmet system, the dataset contains complex large-scale scenes including wild, urban and tunnel, with challenges of drastic motion from wearer and repeated symmetrical geometric structures.
-The datasets used include Residence_5.2, Park_3.2, Subway_3.3, Mall_6.3.
-  • BotanicGarden [31]: Mounted on all-terrain wheeled robot, this dataset consists of diverse natural scenes including riversides, narrow trails and bridges.
-The datasets used include the Garden series.
-  • Handheld datasets: We collected multiple datasets of different scenarios and motion conditions, covering indoor and outdoor scenes like library, campus road and UAV airfield.
+IV-A Experiment Dataset
+The datasets cover a wide range of structured environments (e.g., campus road and garage) and unstructured settings (e.g., botanical gardens), as well as different carriage modes (helmet, vehicle, handheld) to boost the challenge and completeness of assessment.
+  • WHU-Helmet [40]: Collected on compact helmet system with Livox LiDAR, the dataset contains complex large-scale scenes including wild, urban and tunnel, with challenges of drastic motion from wearer and repeated symmetrical geometric structures.
+  • BotanicGarden [41]: Mounted on all-terrain wheeled robot, this dataset consists of diverse natural scenes including riversides, narrow trails and bridges with both Livox Avia and VLP16 in sequence.
+  • Handheld datasets: We collected multiple datasets of different scenarios and motion conditions, covering indoor and outdoor scenes like library, campus road and UAV airfield with Livox Avia.
 The handheld data collection device is shown in Fig.
 4.
-The datasets used include Indoor and Outdoor series.
-In the evaluation of WHU-Helmet, Point-LIO diverged midway in the Residence_5.2 and Subway_3.3 due to severe motion combined with locally degraded scenarios, and diverged in Park_3.2 due to extensive occlusion by pedestrians.
-And in Mall_6.3, the system experienced significant drift during turns due to the reflection from glass walls of the shopping mall.
-In the first three datasets, our AC-LIO framework effectively resists the challenging scenarios mentioned above and outperformed the FAST-LIO in long-term, large-scale localization accuracy.
-Only in the last dataset of WHU-Helmet slightly lower than FAST-LIO.
-For the BotanicGarden datasets, since the sensors are mounted on a wheeled robotic platform, the overall motion is smooth and suitable for evaluating the long-duration drift under regular motion conditions.
-Across the three LIO-rosbags provided by BotanicGarden, the AC-LIO demonstrates higher accuracy, exhibiting better convergence in long-term localization.
-Next is the evaluation of handheld datasets.
-In Indoor_1 UAV airfield and Indoor_3 library scenarios, we conduct intense turning in narrow environment.
-And in the Indoor_2 garage scene, with smooth motion and highly structured surroundings, Point-LIO achieves higher closed-loop accuracy.
-Then in the Outdoor sequences, we handhold the device and attempt different motion conditions including running and sharp turns in campus road scenes.
-Concrete experiment results are demonstrated in Table I, Table II and Fig.
-5 .
-The results indicate the AC-LIO framework generally outperforms the FAST-LIO2 and Point-LIO in terms of accuracy during long-term, large-scale scenes, and exhibits high robustness under challenging scenarios.
+IV-B Odometry Accuracy Evaluation
 Figure 5: Accuracy evaluation in campus road scene.
-The figure shows the result of Outdoor_1.
+The figure shows the result of sequence Outdoor_1.
+Within the dashed box are, in order, the car, the streetlight, and the pattern on the ground.
 We keep running after initialization to enhance the motion intensity and complete a closed loop about 300m.
 The closure trajectory and the map consistency suggest that our AC-LIO achieves better performance.
-IV-B Convergence Criteria Validity Evaluation
-To evaluate the validity of the proposed convergence criteria, we examine the odometry accuracy at different $\eta$ threshold levels.
-While keeping other parameters the same, we only adjust the threshold of convergence criteria.
-Due to the factors such as dynamic objects, cumulative pointcloud map errors, etc., we fine-tune the threshold $\eta$ within the interval of $[\sigma_{t},2\sigma_{t}]$.
-The specific experimental results are shown in Table III.
-  • $\eta=0$: No backpropagation and asymptotic compensation is performed at this situation, reflecting the basic accuracy of the odometry.
-  • $\eta=0.5\sigma_{t}$: Lower than the expected mean of pointcloud residuals, rarely perform backpropagation. (Only considering LiDAR ranging errors, $\mathbb{E}(\mathbf{z})=\sigma_{t}$)
-  • $\eta\in[\sigma_{t},2\sigma_{t}]$: Recommended interval in consideration of other error factors, assuring the backpropagation and asymptotic compensation will be performed with sufficient convergence in previous frame.
-  • $\eta=3\sigma_{t}$: Too high a threshold may lead to backpropagation in case of large initial errors, potentially damaging the system’s accuracy and robustness.
-The experimental results show that, in most cases, selecting the threshold $\eta$ within the range of $[\sigma_{t},2\sigma_{t}]$ achieved higher accuracy.
-Overall, as the threshold increases from zero, the odometry accuracy initially improves because more sufficiently converged states are allowed to perform backpropagation and asymptotic compensation.
-However, as the threshold increases further, states with larger initial errors are also allowed to backpropagate, which reduces the effectiveness of the asymptotic compensation and negatively impacts odometry accuracy.
-TABLE III: Convergence Criteria Validity Evaluation - RMSE of ATE ($m$)
+In odometry accuracy experiments, we compare AC-LIO framework with current state-of-the-art LIO systems including FAST-LIO2, DLIO, SLICT and Point-LIO.
+Some other frameworks are designed for spinning LiDAR hence they’re not included in comparison.
+All algorithms are implemented in C++ and employed on an i9 CPU computer with the ROS in Ubuntu.
+These methods only utilize CPU for computation without the assistance of GPU parallel computing.
+For comparison, we try to fine-tune other methods’ parameters to achieve their best performance.
+Besides, we maintain the threshold of convergence criteria in the extended interval, i.e., $\eta\in[\sigma_{t},2\sigma_{t}]$.
+In the evaluation of WHU-Helmet, the sequences Residence_5.2 and Subway_3.3 introduce aggressive motion combined with locally degraded scenarios, while Park_3.2 and Mall_6.3 suffer from extensive occlusion by pedestrians.
+These factors impose higher demands on the robustness of the LIO system.
+For the BotanicGarden datasets, we test both the solid-state (Livox AVIA) and the spinning (VLP16) LiDAR.
+Since the sensors are mounted on a wheeled robotic platform, the overall motion is smooth and suitable for evaluating long-duration drift under regular motion conditions.
+However, the abundance of unstructured scenes in the garden (e.g., thick woods, bushes) poses challenges for constructing residual constraints.
+In our self-collected datasets, we handheld the sensor suite and completed multiple closed-loop paths in different scenarios.
+In Indoor_1 UAV airfield and Indoor_3 library scenarios, we conduct intense turning in narrow environment.
+For the Outdoor sequences, we handhold the device and attempt different motion conditions including running and sharp turns in campus road scenes.
+Concrete experiment results are demonstrated in Table II and Fig.
+5.
+Our approach further optimizes the residual motion distortion via feedback mechanism based on the convergence criteria, thus improving the accuracy while ensuring the robustness of the system under challenging scenarios.
+The results indicate that our AC-LIO method generally outperforms other benchmark frameworks in terms of accuracy during long-term, large-scale scenes, with the average RMSE among sequences decreased by 30.4% compared to the second best, and exhibits high robustness under challenging scenarios.
+TABLE III: Time Consumption Evaluation ($ms$)
+Ours FAST-LIO2 DLIO SLICT Point-LIO
+Min 4.65 3.15 3.43 21.33 11.95
+Max 74.46 28.48 27.01 345.97 62.36
+Mean 14.62 16.09 9.12 61.97 34.31
+Median 11.71 16.68 8.50 51.49 34.98
+Figure 6: The result of time efficiency evaluation.
+The cross line inside box denotes the median process time.
+Our AC-LIO achieves the second highest efficiency, only behind the lightweight DLIO framework.
+And the maximum iteration time also meets the real-time requirements.
+IV-C Time Efficiency Evaluation
+We calculate the runtime of each algorithm during iteration, with the results shown in Table III and Fig.
+6.
+In terms of average and median processing time, the time efficiency of our AC-LIO framework ranks second place among the tested frameworks, just following the lightweight DLIO framework, and the maximum time for iteration also satisfies realtime requirement (below 100ms for 10Hz LiDAR input).
+Due to the convergence criteria strategy, our method can avoid incurring extra optimization costs in unreliable or unnecessary conditions, thus achieving the balance between system convergence and high time efficiency.
+IV-D Convergence Criteria Validity Evaluation
+TABLE IV: Convergence Criteria Validity Evaluation - RMSE of ATE ($m$)
 00 0 $0.5\sigma_{t}$ $[\sigma_{t},2\sigma_{t}]$ $3\sigma_{t}$
@@
 1.197 1.199 1.152 1.178
-V CONCLUSIONS
-In this paper, we propose AC-LIO, a novel asymptotically and consistently converging LIO framework.
-We reconsider the factors that may introduce errors in long-duration, large-scale scenarios, aiming to further enhance the convergence of the odometry.
-First, during iterative state estimation, we backpropagate the update term along the prior state chain and asymptotically compensate for residual distortion before the next iteration.
-Second, considering the weak correlation between the initial error and motion distortion in the current frame, we propose a convergence criteria based on pointcloud constraints to control backpropagation.
-In summary, the approach of guiding asymptotic compensation based on convergence criteria is designed to promote the consistent convergence of pointcloud registration, ensuring robustness while further improving the odometry accuracy.
-We conducted a series of experiments on public and self-collected datasets, covering various scenarios and motion conditions across multiple scales.
-The experiments show that, compared to other state-of-the-art frameworks, our AC-LIO framework effectively promotes consistent convergence in state estimation and further enhances the accuracy of long-term, large-scale localization and mapping.
+To evaluate the validity of the proposed convergence criteria, we examine the odometry accuracy at different $\eta$ threshold levels.
+While keeping other parameters the same, we only adjust the threshold of convergence criteria.
+Due to the factors such as dynamic objects, cumulative pointcloud map errors, etc., we fine-tune the threshold $\eta$ within the interval of $[\sigma_{t},2\sigma_{t}]$.
+The specific experimental results are shown in Table IV.
+  • $\eta=0$: No backpropagation and asymptotic compensation is performed at this situation, reflecting the basic accuracy of the odometry.
+  • $\eta=0.5\sigma_{t}$: Lower than the expected mean of pointcloud residuals, rarely perform backpropagation. (Only considering LiDAR ranging errors, $\mathbb{E}({z})=\sigma_{t}$)
+  • $\eta\in[\sigma_{t},2\sigma_{t}]$: Recommended interval in consideration of other error factors, assuring the backpropagation and asymptotic compensation will be performed with sufficient convergence in previous frame.
+  • $\eta=3\sigma_{t}$: Too high a threshold may lead to backpropagation in case of large initial errors, potentially damaging the system’s accuracy and robustness.
+The experimental results show that, in most cases, selecting the threshold $\eta$ within the range of $[\sigma_{t},2\sigma_{t}]$ achieved higher accuracy.
+Overall, as the threshold increases from zero, the odometry accuracy initially improves because more sufficiently converged states are allowed to perform backpropagation and asymptotic compensation.
+However, as the threshold increases further, states with larger initial errors are also allowed to backpropagate, which reduces the effectiveness of the asymptotic compensation and negatively impacts odometry accuracy.
+V Conclusion and Future Work
+In this paper, we proposed AC-LIO, a novel asymptotically and consistently converging LIO framework.
+We reconsider the factors that may introduce errors in long-duration, large-scale scenarios, aiming to further enhance the convergence of LIO system.
+The approach of guiding asymptotic compensation based on convergence criteria is designed to promote consistent convergence of pointcloud registration, further improving the odometry accuracy while ensuring system robustness.
+Numerous experiment results indicate that our framework’s average RMSE is reduced by about 30.4% over the second best result, achieving higher accuracy in long-term, large-scale localization and mapping.
+Future Work: We notice the performance of back propagation depends on the guidance of convergence criteria to some extend.
+The introduction of convergence criteria is essentially an attempt to overcome the disruption of the initial error on the back propagation during the iteration update.
+In future work, we will continue to study the convergence assessment of odometry, the combination of asymptotic backpropagation with other strategies, and the perception and adaptation of degenerate scenarios, to further explore the error-analyzable and scenario-interpretable SLAM framework.

### v3

@@
-AC-LIO: Towards Asymptotic and Consistent Convergence in LiDAR-Inertial Odometry
+AC-LIO: Towards Asymptotic Compensation for Distortion in LiDAR-Inertial Odometry via Selective Intra-Frame Smoothing
 Abstract
-Existing LiDAR-Inertial Odometry (LIO) methods typically utilize the prior state trajectory derived from the IMU integration to compensate for the motion distortion within LiDAR frames.
-However, discrepancies between the prior and actual trajectory can lead to residual distortions that compromise the consistency of the LiDAR frame with its corresponding geometric environment.
+Existing LiDAR-Inertial Odometry (LIO) methods typically utilize the prior trajectory derived from the IMU integration to compensate for the motion distortion within LiDAR frames.
+However, discrepancies between the prior and true trajectory can lead to residual motion distortions that compromise the consistency of LiDAR frame with its corresponding geometric environment.
 This imbalance may result in pointcloud registration becoming trapped in local optima, thereby exacerbating drift during long-term and large-scale localization.
-To address the issue, we propose a novel asymptotically and consistently converging LIO framework dubbed AC-LIO.
-Our key idea is to back propagate current update term based on the prior state chain, and asymptotically compensate for the residual distortion during iteration.
-Moreover, considering the weak correlation between previous error and current distortion, we establish convergence criteria based on the pointcloud constraints to regulate the backpropagation.
-This method of guiding asymptotic distortion compensation using convergence criteria subtly enhances the consistent convergence of pointcloud registration, futher improving the accuracy and robustness of LIO system.
-Extensive experiments demonstrate that our AC-LIO framework significantly promotes consistent convergence in state estimation compared to prior arts, with about 30.4% reduction in average RMSE over the second best result, leading to marked improvements in the accuracy of long-term and large-scale localization and mapping.
+To this end, we propose a novel LIO framework with selective intra-frame smoothing dubbed AC-LIO.
+Our core idea is to asymptotically backpropagate current update term and compensate for residual motion distortion under the guidance of convergence criteria, aiming to improve the accuracy of discrete-state LIO system with minimal computational increase.
+Extensive experiments demonstrate that our AC-LIO framework further enhances odometry accuracy compared to prior arts, with about 30.4% reduction in average RMSE over the second best result, leading to marked improvements in the accuracy of long-term and large-scale localization and mapping.
+Our code and demo are available on https://cyberkona.github.io/publication/ac-lio/.
+Index Terms:
+LiDAR-inertial odometry (LIO), simultaneous localization and mapping (SLAM), asymptotic distortion compensation, convergence criteria, selective intra-frame smoothing.
 I Introduction
-With the advancements in Simultaneous Localization and Mapping (SLAM) technology, robots are increasingly achieving precise perception of their external environments and their own motion[1, 2, 3].
-As a vital subset of SLAM, LiDAR-Inertial Odometry (LIO) has gained widespread application in autonomous driving [4], unmanned logistics [5], humanoid robot [6] and other fields, for its high robustness and accuracy [7, 8].
-Compared to Visual-Inertial Odometry, the LIO system can directly obtain depth observations of surroundings [9], without the need to estimate the depth of feature points [10, 11], and exhibits better adaptability to different lighting conditions [12, 13].
+With the advancements in Simultaneous Localization and Mapping (SLAM) technology, robots are increasingly achieving precise perception of external environments and their own motion[1, 2, 3].
+As a vital subset of SLAM, LiDAR-Inertial Odometry (LIO) has gained widespread application in autonomous driving [4], unmanned logistics [5] and other fields, for its high robustness and accuracy [6, 7].
+Compared to Visual-Inertial Odometry, LIO system can directly obtain depth observations of surroundings, without the need to estimate the depth of feature points [8], and exhibits better adaptability to different lighting conditions [9, 10].
 Figure 1: (a) A principal comparison of our AC-LIO with traditional approaches.
-Conventional LIO typically considers an initial distortion compensation via IMU, yet the residual distortion may prevent the registration from further convergence.
-In contrast, our AC-LIO employs an asymptotic backpropagation to promote consistent convergence in state estimation.
+Conventional LIO typically considers an initial distortion compensation via IMU, yet the residual motion distortion may prevent the registration from further convergence.
+Our AC-LIO employs a selective intra-frame smoothing strategy to asymptotically backpropagate and compensate for residual distortion.
 (b) Large scale localization and mapping result of BotanicGarden dataset.
-We achieve better accuracy performance than other benchmark frameworks on its sequence 1008_01 with VLP16 LiDAR.
+We achieve better accuracy performance than other benchmark frameworks on its sequence Garden_1008 with VLP16 LiDAR.
 However, as LIO system continues to evolve, there still remain several fundamental issues to be addressed [11, 12, 13].
@@
 1, traditional methods [15, 16, 17] typically only consider using IMU integration for the initial compensation of motion distortion, without further optimization of the residual distortion in the state estimation.
-This residual distortion will compromise the consistency between the LiDAR frame with its corresponding geometric environment, and it grows with increasing motion non-linearity.
+This residual motion distortion will compromise the consistency between the LiDAR frame with its corresponding geometric environment, and it grows with increasing motion non-linearity.
 Therefore, in conventional LIO systems, such residual error may suppress further convergence of pointcloud registration during the iterative update, ultimately impacting the odometry accuracy and causing severe drift.
-In this letter, we propose AC-LIO, a novel asymptotically and consistently converging LIO framework.
-Our AC-LIO enjoys two key technical contributions.
-First, during the iterative state estimation, we introduce an asymptotic distortion compensation (Sec.
-III-C) method that backwards propagate the current update term based on the prior state chain, and asymptotically compensate the residual distortion before next iteration .
-This approach could enhance the consistency between the current LiDAR frame and the represented geometric environment, and avoid the pointcloud registration from falling into local optimum due to the residual distortion, thereby promoting the consistent convergence of pointcloud registration and improving the accuracy of long-term and large-scene localization.
-Second, considering the weak correlation between initial error and motion distortion of current frame, we also propose the convergence criteria (Sec.
+Besides, we notice that few works directly apply the reliable and concise Rauch-Tung-Striebel (RTS) smoother for continuous optimization of motion distortion within discrete state LIO frameworks.
+While smoothing strategies have been widely adopted in continuous state estimation and several sliding window factor graph frameworks, they primarily focus on joint optimization across multiple states, which often leads to substantial computational consumption.
+However, since the LIO systems pursue a balance between real-time performance and accuracy, we seek a solution tailored for discrete state LIO frameworks that, without introducing additional estimates or state variables, leverage smoothing methods to achieve asymptotical compensation for residual motion distortion during iterative state estimation.
+Moreover, some redundant computation can be avoid as the state covariance chains are available from IMU prior propagation.
+In this work, we propose AC-LIO, a novel discrete state LIO framework with selective intra-frame smoothing.
+Our AC-LIO introduces two key technical contributions.
+First, during iterative state estimation, we propose an asymptotic distortion compensation (Sec.
+III-C) method that backwards propagate the current update term based on the prior state chain, and asymptotically compensate fr the residual distortion before next iteration.
+This approach could enhance the consistency between the current LiDAR frame and the represented geometric environment, and avoid the pointcloud registration from falling into local optimum due to the residual distortion, thereby promoting the accuracy of long-term and large-scene localization.
+Second, due to the lack of correlation between initial error and motion distortion of current frame, we also propose the convergence criteria (Sec.
 III-D) based on the point-to-plane residuals to control the back propagation.
-The criteria is used to assess the convergence of the pointcloud registration and ensure that the backpropagation and asymptotic compensation only happens under the condition of sufficient convergence in previous LiDAR frame, hence guaranteeing the robustness of LIO system.
+We apply the criteria to assess the convergence of the pointcloud registration and ensure that the backpropagation and asymptotic compensation only happens under the condition of sufficient convergence in previous LiDAR frame, hence guaranteeing the robustness of LIO system.
 Compared to other strategies such as continuous-time estimation or frame segmentation, our method leverages this criteria to construct a feedback mechanism that promotes state convergence without additional estimation, achieving a balance between localization accuracy and computational efficiency (Sec.
-IV-C).
-It can be seamlessly embedded into the iterative Kalman Filter framework [21], offering great flexibility and scalability.
+IV-E).
+It can be seamlessly embedded into the iterative Kalman Filter framework, offering great flexibility and scalability.
 In summary, our main contributions are as follows:
-  • We propose a novel framework AC-LIO with asymptotic distortion compensation that showing plausible consistent convergence in long-term and large-scale scenarios.
+  • We propose a novel framework AC-LIO with asymptotic distortion compensation to further suppress residual distortion during iterative update, promoting system accuracy in long-term and large-scale scenarios.
   • We introduce the convergence criteria based on point-to-plane residuals to evaluate the convergence and guide backpropagation.
@@
 The results show that our method achieves higher accuracy compared to current state-of-the-art frameworks.
+We also release our code to support and contribute to the open-source community.
 II RELATED WORK
@@
 However, this assumption only works effectively when LiDAR frame interval is enough short and the motion is slow and smooth.
-In case of drastic velocity variation, the constant velocity model cannot accurately characterize the current motion state, and this will result in increased drift and even odometry divergence.
+In case of drastic velocity variation, constant velocity model cannot accurately characterize current motion state, leading to increased drift and even odometry divergence.
 CT-ICP [20] optimize the start and end states of LiDAR frame, and construct a continuous time trajectory by interpolation.
 But the real movement may not follow the interpolated trajectory, as well as the logical constraint across adjacent frames is hard to determine.
-I2EKF-LO [25] utilizes extra prediction iteration to estimate motion distortion, yet the update of the process noise still depends on the hyperparameters.
-Traj-LO [26] leverages the high frequency of LiDAR points and employs more linear segments within frame.
+I2EKF-LO [21] utilizes extra prediction iteration to estimate motion distortion, yet the update of process noise still depends on some hyperparameters.
+Traj-LO [22] leverages high frequency of LiDAR points and employs more linear segments within frame.
 This approach allows more accurate trajectory characterization and suppression of motion aberrations during high speed motion, but also introduces more computational cost.
@@
 Continuous State LiDAR-Inertial Odometry.
-To characterize trajectorie more precisely in state estimation, many frameworks come up with new optimization solutions.
-CLINS [33] introduces the B-spline [34] for continuous time state estimation, while SLICT [35] employs continuous-time factors to represent LiDAR constraints within a sliding window.
-These frameworks further optimize motion distortion with continuous-time model.
-However, the unconstrained trajectory would not always conform to the predefined model due to time resolution limitations in high-speed, dynamic motion.
+To characterize trajectories more precisely in state estimation, many frameworks come up with new optimization solutions.
+CLINS [29] introduces the B-spline for continuous state estimation, while SLICT [30] employs continuous-time factors to represent LiDAR constraints within a sliding window.
+However, the unconstrained trajectory would not always conform to predefined model due to time resolution limitations in high-speed, dynamic motion.
 Moreover, the continuous-time optimization strategy may introduce a higher computational cost.
-DLIO [36] uses analytical equations to construct a continuous trajectory for precise motion correction, after coarse deskewing through IMU integration.
-But the method remains limited by IMU’s inherent accuracy and neglects further undistortion during the registration process.
-In contrast, our proposed AC-LIO framework utilizes the convergence criteria to guide the backpropagation process, to achieve the asymptotic compensation of residual motion distortion.
-As a result, our approach can promote the consistent convergence of pointcloud registration during iterative update, further improving the accuracy over long-term and large-scale scenarios while ensuring the system robustness.
+DLIO [31] uses analytical equations to model a continuous trajectory for precise motion correction, after coarse deskewing with IMU integration.
+Yet it remains limited by IMU’s inherent accuracy and neglects further undistortion during registration process.
+In contrast, our proposed AC-LIO framework utilizes the convergence criteria to guide the backpropagation and asymptotic compensation for residual motion distortion, further improving the accuracy over long-term and large-scale scenarios while ensuring the system robustness.
 III METHODOLOGY
+III-A System Overview
+Figure 2: (a) The system overview of our AC-LIO framework.
+During iterative ESKF, the back propagation and asymptotic compensation for pointcloud distortion are conducted based on the convergence criteria.
+(b) The on-manifold propagation.
+The blue shows the forward propagation of IMU integration.
+The green indicates the backward recursion of the update term according to the propagation matrix and noise relation of error state chains.
+The system overview of our AC-LIO framework is shown in Fig.
+2, and the important notations is shown in Table I.
+First, the IMU observations are fed into the forward propagation module to obtain prior trajectories.
+Then the LiDAR data utilize the prior state chain for initial compensation and then sent to the Error State Kalman Filter (ESKF) for state estimation [32], where the average point-to-plane residual (APR) is calculated after each update.
+During the iteration, if the APR of previous LiDAR frame is lower than the convergence criteria and the current APR exceeds it, then the backpropagation of update residual will be executed.
+The LiDAR frame will be re-compensated to promote the convergence of registration after each backpropagation.
+Otherwise, if the conditions are not met, the backpropagation will be skipped and iteration continues.
+Once reaching the maximum of iterations, the system will merge the LiDAR frame under the posterior pose into pointcloud map and publish the current odometry.
+III-B ESKF State Estimation
 TABLE I: Important Notations
@@
 $\bar{\mathbf{x}},\check{\mathbf{x}}$ The updated state $\mathbf{x}$ with or without back propagation.
-$\mathbf{F_{a}}^{i|i+n}$ The propagation matrix from state $i$ to $i+n$.
+$\mathbf{F_{p}}^{i|i+n}$ The propagation matrix from state $i$ to $i+n$.
 $\mathbf{G}_{i}^{i+n}$ The backward gain matrix of state $i$ with respect to $i+n$.
@@
 $\sigma_{t}$ The expectation of average point-to-plane residual.
-III-A System Overview
-Figure 2: (a) The system overview of our AC-LIO framework.
-During iterative ESKF, asymptotic compensation of pointcloud distortion is conducted based on the convergence criteria, to promote consistent convergence of state estimation.
-(b) The on-manifold propagation.
-The blue shows the forward propagation of IMU integration.
-The green indicates the backward recursion of the update term according to the propagation matrix and noise relation.
-The system overview of our AC-LIO framework is shown in Fig.
-2, and the important notations is shown in Table I.
-First, the IMU observations are fed into the forward propagation module to obtain prior trajectories.
-Then the LiDAR data utilize the prior state chain for initial compensation and then sent to the Error State Kalman Filter (ESKF) for state estimation [37], where the average point-to-plane residual (APR) is calculated after each update.
-During the iteration, if the APR of previous LiDAR frame is lower than the convergence criteria and the current APR exceeds it, then the backpropagation of update residual will be executed.
-The LiDAR frame will be re-compensated to promote the convergence of registration after each backpropagation.
-Otherwise, if the conditions are not met, the backpropagation will be skipped and iteration continues.
-Once reaching the maximum of iterations, the system will merge the LiDAR frame under the posterior pose into pointcloud map and publish the current odometry.
-III-B ESKF State Estimation
 For efficiency and accuracy concerns, the AC-LIO we proposed utilizes the ESKF framework for state estimation.
@@
 […]
-$\begin{gathered}\quad\quad\delta\mathbf{x}_{i}\doteq\begin{bmatrix}\delta{\bm{\theta}}&\delta\mathbf{p}_{I}^{G}&\delta\mathbf{v}_{I}&\delta\mathbf{b}_{\bm{\omega}}&\delta\mathbf{b}_{\mathbf{a}}&\delta\mathbf{g}^{G}\end{bmatrix}^{T}\\ \quad\delta\hat{\mathbf{x}}_{i+1}=\mathbf{F_{a}}\delta\check{\mathbf{x}}_{i}+\mathbf{F_{n}}\mathbf{n}_{i}\\ \quad\hat{\mathbf{P}}_{i+1}=\mathbf{F_{a}}\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}+\mathbf{F_{n}}\mathbf{Q}\mathbf{F_{n}}^{T}\end{gathered}$ (2)
+$\begin{gathered}\quad\quad\delta\mathbf{x}_{i}\doteq\begin{bmatrix}\delta{\bm{\theta}}&\delta\mathbf{p}_{I}^{G}&\delta\mathbf{v}_{I}&\delta\mathbf{b}_{\bm{\omega}}&\delta\mathbf{b}_{\mathbf{a}}&\delta\mathbf{g}^{G}\end{bmatrix}^{T}\\ \quad\delta\hat{\mathbf{x}}_{i+1}=\mathbf{F_{p}}\delta\check{\mathbf{x}}_{i}+\mathbf{F_{n}}\mathbf{n}_{i}\\ \quad\hat{\mathbf{P}}_{i+1}=\mathbf{F_{p}}\check{\mathbf{P}}_{i}\mathbf{F_{p}}^{T}+\mathbf{F_{n}}\mathbf{Q}\mathbf{F_{n}}^{T}\end{gathered}$ (2)
 Assume the extrinsic $\mathbf{T}_{L}^{I}$ between LiDAR and IMU is already known.
@@
 $\begin{gathered}\mathbf{z}=\mathbf{u}(\hat{\mathbf{p}}^{G}-\mathbf{q}^{G})^{T}\\ \mathbf{H}=\begin{bmatrix}-\mathbf{u}\hat{\mathbf{R}}_{I}^{G}(\mathbf{T}_{L}^{I}\mathbf{p}^{L})_{\wedge}&\mathbf{u}&\mathbf{0}_{3\times 12}\end{bmatrix}\end{gathered}$ (3)
-Considering the nonlinear errors in the $\mathbf{F_{a}}$, $\mathbf{F_{n}}$ and $\mathbf{H}$, the estimation may require iteration to promote convergence.
+Considering the nonlinear errors in the $\mathbf{F_{p}}$, $\mathbf{F_{n}}$ and $\mathbf{H}$, the estimation may require iteration to promote convergence.
 Let $\mathbf{A}$ denote the Jacobian of update component with respect to $\delta\mathbf{x}$, and according to [15], the update equation is:
@@
 $\begin{gathered}\check{\mathbf{x}}_{i+1}\Leftarrow\hat{\mathbf{x}}_{i+1}^{j+1},\;\check{\mathbf{P}}_{i+1}\Leftarrow(\mathbf{I}-\mathbf{K}\mathbf{H})\hat{\mathbf{P}}_{i+1}\end{gathered}$ (5)
-With the iterative update of ESKF, our AC-LIO could further promote the consistent convergence of the state estimation via the asymptotic distortion compensation and convergence criteria introduced in following sections.
+With the iterative update of ESKF, our AC-LIO could further promote the convergence of the state estimation via the asymptotic distortion compensation and convergence criteria introduced in following sections.
 III-C Asymptotic Distortion Compensation
@@
 With reference to the Rauch-Tung-Striebel fixed interval smooth strategy [34], we can back propagate the update term $\delta\mathbf{x}$ based on the prior state chain.
-The contribution of previous state in current update term just corresponds to the proportion of propagation covariance $\mathbf{F_{a}}\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}$ within current covariance $\hat{\mathbf{P}}_{i+1}$.
+The contribution of previous state in current update term just corresponds to the proportion of propagation covariance $\mathbf{F_{p}}\check{\mathbf{P}}_{i}\mathbf{F_{p}}^{T}$ within current covariance $\hat{\mathbf{P}}_{i+1}$.
 Then, the gain matrix $\mathbf{G}$ and the backward recursion of $\delta\mathbf{x}$ can be denoted as follows:
-$\begin{gathered}\mathbf{G}=\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}\hat{\mathbf{P}}_{i+1}^{-1}\\ \delta\hat{\mathbf{x}}_{i}^{j}=\mathbf{G}\cdot\delta\hat{\mathbf{x}}_{i+1}^{j}\end{gathered}$ (6)
+$\begin{gathered}\mathbf{G}=\check{\mathbf{P}}_{i}\mathbf{F_{p}}^{T}\hat{\mathbf{P}}_{i+1}^{-1}\\ \delta\hat{\mathbf{x}}_{i}^{j}=\mathbf{G}\cdot\delta\hat{\mathbf{x}}_{i+1}^{j}\end{gathered}$ (6)
 Furthermore, this sequential backpropagation method can be extend to any states within the LiDAR frame.
 It allows us to selectively update several anchor in the prior state chain, thus reducing the computational complexity of backpropagation and distortion compensation.
-As for any states $\mathbf{x_{i}}$ and $\mathbf{x_{i+n}}$ in the prior state chain, the propagation matrix and backward gain matrix can be noted as $\mathbf{F_{a}}^{i|i+n}$ and $\mathbf{G}_{i}^{i+n}$, then the backpropagation relationship should be:
-$\begin{gathered}\mathbf{F_{a}}^{i|i+n}=\prod_{k=i}^{i+n-1}\mathbf{F_{a}}^{k|k+1}\\ \mathbf{G}_{i}^{i+n}=\check{\mathbf{P}}_{i}(\mathbf{F_{a}}^{i|i+n})^{T}(\hat{\mathbf{P}}_{i+n})^{-1}\\ \check{\mathbf{x}}_{i}^{j}=\check{\mathbf{x}}_{i}\boxplus(\mathbf{G}_{i}^{i+n}\cdot\delta\hat{\mathbf{x}}_{i+n}^{j})\end{gathered}$ (7)
-After each backpropagation of the update term $\hat{\mathbf{x}}_{i+n}^{j}$, the residual pointcloud distortion can continue to be compensated based on the smoothed trajectory.
+As for any states $\mathbf{x_{i}}$ and $\mathbf{x_{i+n}}$ in the prior state chain, the propagation matrix and backward gain matrix can be noted as $\mathbf{F_{p}}^{i|i+n}$ and $\mathbf{G}_{i}^{i+n}$, then the backpropagation relationship should be:
+$\begin{gathered}\mathbf{F_{p}}^{i|i+n}=\prod_{k=i}^{i+n-1}\mathbf{F_{p}}^{k|k+1}\\ \mathbf{G}_{i}^{i+n}=\check{\mathbf{P}}_{i}(\mathbf{F_{p}}^{i|i+n})^{T}(\hat{\mathbf{P}}_{i+n})^{-1}\\ \check{\mathbf{x}}_{i}^{j}=\check{\mathbf{x}}_{i}\boxplus(\mathbf{G}_{i}^{i+n}\cdot\delta\hat{\mathbf{x}}_{i+n}^{j})\end{gathered}$ (7)
+After each backpropagation of the update term $\hat{\mathbf{x}}_{i+n}^{j}$, where the $j$ refers the current iteration times, the residual pointcloud distortion can continue to be compensated based on the smoothed trajectory.
 For the point $\mathbf{p}^{L}$ in LiDAR frame, and the state corresponding to its timestamp $t$ after the $j$th backpropagation can be denoted as $\hat{\mathbf{x}}_{t}^{j}$, then the updated coordinates can be noted as:
@@
 III-D Convergence Criteria
+Due to the lack of correlation between the previous error and the current motion distortion within a LiDAR frame, we find it necessary to introduce the convergence criteria to regulate the asymptotic distortion compensation.
+If the previous LiDAR frame has sufficiently converged, the current frame may start with a relatively small initial error, allowing the backpropagation to effectively compensate for motion distortion during iterative state estimation.
+However, if previous LiDAR frame has not converged well, then the initial error may degrade the system accuracy via the backpropagation.
+Therefore, a simple and efficient method is required to detect the convergence status of each LiDAR frame, in order to guide the asymptotic distortion compensation.
 Figure 3: The relationship between original ranging variance of LiDAR point and point-to-plane residual variance, where the residual $z$ is a scalar.
-To exploit the statistical properties of pointcloud, during the iterative state estimation, we select the mean value of the point-to-plane residual (abbreviated as APR above), as the criteria for evaluating the convergence of pointcloud registration.
-For simplification, assuming the pointcloud registration fully converges to the global optimum and only considering the ranging accuracy of LiDAR, we can roughly model the expectation of the APR and provide a threshold reference for the convergence criteria.
+To exploit the statistical properties of pointcloud, during iterative state estimation, we select the mean value of the point-to-plane residual (abbreviated as APR above), as the criteria for evaluating the convergence of pointcloud registration.
+For simplification, assuming the pointcloud registration converges to the true pose and only considering the ranging accuracy of LiDAR, we can roughly model the expectation of the APR and provide a threshold reference for the convergence criteria.
 Denote the incident angle of point $\mathbf{p}$ to the fitted plane $S$ as $\phi$, it can be expressed as $\cos^{-1}\tfrac{|\mathbf{u}\cdot\hat{\mathbf{p}}^{G}|}{|\mathbf{u}|\cdot|\hat{\mathbf{p}}^{G}|}$ according to Eq.
@@
 $\begin{gathered}\mathbb{E}({z})=\frac{2}{\pi}\int_{0}^{\frac{\pi}{2}}\mathbb{E}({z}_{\perp})\cos{\phi}\;d\phi=\frac{2}{\pi}\sigma\end{gathered}$ (11)
-Therefore, assuming the pointcloud registration converges to the global optimum, and only considering the LiDAR ranging error, the expectation of APR should be ${2\sigma}/{\pi}$.
+Therefore, assuming the pointcloud registration converges to the true pose, and only considering the LiDAR ranging error, the expectation of APR should be ${2\sigma}/{\pi}$.
 If the mean value of current point-to-plane residual is lower than this criteria, the last registration can be regarded as sufficiently converged.
@@
 Noting that $\sigma_{t}=\mathbb{E}({z})$ and the selected threshold of convergence criteria as $\eta$, due to the factors such as dynamic objects and cumulative errors in pointcloud map, the threshold $\eta$ can be appropriately extended to the interval of $[\sigma_{t},2\sigma_{t}]$ in practice.
+We collect a sequence of static-movement-static behavior with Livox AVIA, to evaluate the validity of our proposed convergence criteria.
+As shown in Fig.
+8, the value of $\sigma_{t}$ is very close to the APR during the static phases.
 In detail, we perform two conditional judgments in iterative state estimation: First, evaluate the previous LiDAR frame, if its APR is less than the threshold $\eta$, it can be considered to be sufficiently converged and we will continue.
@@
 IV EXPERIMENT RESULTS
+In this section, we qualitatively and quantitatively evaluate AC-LIO framework through a series of experiments with various public and self-collected datasets.
+To examine the performance of our approach in long-term and large-scale scenarios, we conduct extensive experiments on the odometry accuracy, the validity of convergence criteria, and the time efficiency of system.
+IV-A Experiment Dataset
+The datasets cover a wide range of structured environments (e.g., campus road and garage) and unstructured settings (e.g., botanical gardens), as well as different carriage modes (helmet, vehicle, handheld) to boost the challenge and completeness of assessment.
+  • WHU-Helmet [35]: Collected on compact helmet system with Livox LiDAR, the dataset contains complex large-scale scenes including wild, urban and tunnel, with challenges of drastic motion from wearer and repeated symmetrical geometric structures.
+  • BotanicGarden [36]: Mounted on all-terrain wheeled robot, this dataset consists of diverse natural scenes including riversides, narrow trails and bridges with both Livox Avia and VLP16 in sequence.
+  • Handheld datasets: We collected multiple datasets of different scenarios and motion conditions, covering indoor and outdoor scenes like library, campus road and UAV airfield with Livox Avia.
+The handheld data collection device is shown in Fig.
+4.
+Figure 4: Our handheld sensor suite for self-collected dataset, including solid-state LiDAR Livox Avia with built-in BMI088 IMU, and three other cameras.
 TABLE II: Odometry Accuracy Evaluation - RMSE of ATE & End to End Errors($m$)
@@
 WHU1 Helmet Livox AVIA Residence_5.2 (452.9m) 0.415 0.505 4.660 1.550 $\times$
-Park_3.2 (628.9m) 0.582 0.618 7.725 0.602 $\times$
+Parking_3.2 (628.9m) 0.582 0.618 7.725 0.602 $\times$
 Subway_3.3 (827.0m) 1.670 3.752 3.688 1.967 $\times$
 Mall_6.3 (1006.5m) 0.216 0.208 3.238 0.274 15.409
-Botanic1 Garden Livox AVIA 1005_05 (566.8m) 0.773 1.053 0.640 7.931 0.917
-1006_03 (686.2m) 1.796 2.075 3.017 6.527 2.471
-1008_01 (855.9m) 1.152 1.577 14.829 4.445 1.590
-VLP16 1005_05 (566.8m) 0.175 0.244 0.565 37.425 0.750
-1006_03 (686.2m) 0.377 0.388 0.393 $\times$ 0.445
-1008_01 (855.9m) 0.339 0.363 0.694 $\times$ 0.457
+Botanic1 Garden Livox AVIA Garden_1005 (566.8m) 0.773 1.053 0.640 7.931 0.917
+Garden_1006 (686.2m) 1.796 2.075 3.017 6.527 2.471
+Garden_1008 (855.9m) 1.152 1.577 14.829 4.445 1.590
+VLP16 Garden_1005 (566.8m) 0.175 0.244 0.565 37.425 0.750
+Garden_1006 (686.2m) 0.377 0.388 0.393 $\times$ 0.445
+Garden_1008 (855.9m) 0.339 0.363 0.694 $\times$ 0.457
 Handheld2 Livox AVIA Indoor_1 (~100m) 0.893 1.133 $\times$ 1.960 5.401
@@
   • $\times$  denotes the system severely diverged midway, and $\underline{\hskip 7.11317pt}$  denotes the second best result compared to bolded result.
-In this section, we qualitatively and quantitatively evaluate AC-LIO framework through a series of experiments with various public and self-collected datasets.
-To examine the performance of our approach in long-term and large-scale scenarios, we conduct extensive experiments on the odometry accuracy, the time efficiency of system, and the validity of convergence criteria.
-Figure 4: Our handheld sensor suite used in self-collected dataset, including the solid-state LiDAR Livox Avia with the built-in BMI088 IMU, and three other cameras.
-IV-A Experiment Dataset
-The datasets cover a wide range of structured environments (e.g., campus road and garage) and unstructured settings (e.g., botanical gardens), as well as different carriage modes (helmet, vehicle, handheld) to boost the challenge and completeness of assessment.
-  • WHU-Helmet [40]: Collected on compact helmet system with Livox LiDAR, the dataset contains complex large-scale scenes including wild, urban and tunnel, with challenges of drastic motion from wearer and repeated symmetrical geometric structures.
-  • BotanicGarden [41]: Mounted on all-terrain wheeled robot, this dataset consists of diverse natural scenes including riversides, narrow trails and bridges with both Livox Avia and VLP16 in sequence.
-  • Handheld datasets: We collected multiple datasets of different scenarios and motion conditions, covering indoor and outdoor scenes like library, campus road and UAV airfield with Livox Avia.
-The handheld data collection device is shown in Fig.
-4.
-IV-B Odometry Accuracy Evaluation
 Figure 5: Accuracy evaluation in campus road scene.
@@
 The closure trajectory and the map consistency suggest that our AC-LIO achieves better performance.
-In odometry accuracy experiments, we compare AC-LIO framework with current state-of-the-art LIO systems including FAST-LIO2, DLIO, SLICT and Point-LIO.
-Some other frameworks are designed for spinning LiDAR hence they’re not included in comparison.
+IV-B Experiment Settings:
+  • We compared AC-LIO with current sota LIO frameworks that employ various methods to mitigate motion distortion, rather than LO frameworks without IMU integration.
+Some frameworks are designed for spinning LiDAR hence they are not included in comparison.
+Among the LIO frameworks, FAST-LIO2 compensates for distortion via IMU prior trajectory before iterative estimation, DLIO applies a coarse-to-fine pointcloud deskewing approach, SLICT utilizes continuous-time state estimation to optimize the motion distortion, and Point-LIO is an extremely high frequency LIO with a point-by-point strategy to avoid motion distortion.
   • All algorithms are implemented in C++ and employed on an i9 CPU computer with the ROS in Ubuntu.
 These methods only utilize CPU for computation without the assistance of GPU parallel computing.
-For comparison, we try to fine-tune other methods’ parameters to achieve their best performance.
-Besides, we maintain the threshold of convergence criteria in the extended interval, i.e., $\eta\in[\sigma_{t},2\sigma_{t}]$.
-In the evaluation of WHU-Helmet, the sequences Residence_5.2 and Subway_3.3 introduce aggressive motion combined with locally degraded scenarios, while Park_3.2 and Mall_6.3 suffer from extensive occlusion by pedestrians.
+As our AC-LIO is implemented on the basis of FAST-LIO2, we control the same parameters especially with FAST-LIO2 (e.g., imu_noise, voxel_size, max_iteration, point_filter_num) for fair comparison.
+And we make every effort to fine-tune the parameters of other methods to achieve their best performance.
+  • Since our AC-LIO computes the convergence criteria to guide the asymptotic distortion compensation during iterative update, we fix the iteration times for every epoch (i.e., single LiDAR frame).
+In contrast, FAST-LIO2 jumps out the iteration as the update gradient belows a threshold, which may lead to an accuracy difference with our AC-LIO.
+Therefore, we conducted ablation studies specifically (Sec.
+IV-D) to validate the performance of selective intra-frame smoothing.
+IV-C Odometry Accuracy Evaluation
+Figure 6: Accuracy evaluation in Botanic Garden dataset.
+The figure shows the result of sequence (a) Garden_1006 and (b) Garden_1008.
+For each sequence, the left part presents the trajectory comparison, with the top 10% states aligned to the ground truth, and the right part displays the cumulative error curve of APE (m) along with the corresponding real-time process efficiency.
+The results suggest that our AC-LIO achieves lower cumulative error while maintaining relatively good time efficiency (Sec.
+IV-E).
+The average runtime of our method is only slightly higher than the FAST-LIO2 framework.
+The odometry accuracy evaluation aims to demostrate the overall accuracy performance of selected LIO frameworks and our AC-LIO.
+In the evaluation of WHU-Helmet, the sequences Residence_5.2 and Subway_3.3 introduce aggressive motion combined with locally degraded scenarios, while Parking_3.2 and Mall_6.3 suffer from extensive occlusion by pedestrians.
 These factors impose higher demands on the robustness of the LIO system.
@@
 For the Outdoor sequences, we handhold the device and attempt different motion conditions including running and sharp turns in campus road scenes.
-Concrete experiment results are demonstrated in Table II and Fig.
-5.
-Our approach further optimizes the residual motion distortion via feedback mechanism based on the convergence criteria, thus improving the accuracy while ensuring the robustness of the system under challenging scenarios.
+Concrete experiment results are demonstrated in Table II, Fig.
+5 and Fig.
+6.
+Our approach further optimizes the residual motion distortion via selective intra-frame smoothing strategy, thus improving the accuracy while ensuring the robustness of the system under challenging scenarios.
 The results indicate that our AC-LIO method generally outperforms other benchmark frameworks in terms of accuracy during long-term, large-scale scenes, with the average RMSE among sequences decreased by 30.4% compared to the second best, and exhibits high robustness under challenging scenarios.
-TABLE III: Time Consumption Evaluation ($ms$)
+IV-D Convergence Criteria Evaluation
+TABLE III: Convergence Criteria Validity Evaluation - RMSE of ATE ($m$)
+[…]
+Garden_1005 (566.8m) AVIA 1.053 1.063 1.063 1.004 0.943 0.743 0.907 1.120
+VLP16 0.244 0.244 0.241 0.226 0.175 0.193 0.208 0.233
+Figure 7: Validity evaluation of convergence criteria.
+These two figures present the ablation experiments on convergence criteria using Garden_1005 sequence: (a) for Livox AVIA and (b) for VLP16, showing the generalization between different LiDAR sensors.
+The gray dashed line refers the accuracy of FAST-LIO2.
+As threshold $\eta$ increases from zero, the accuracy improves because more sufficiently converged states are allowed to perform backpropagation.
+However, as $\eta$ continues to increase, states with larger initial errors are also allowed to backpropagate, which gradually degrades the overall accuracy.
+To evaluate the rationality and validity of the proposed convergence criteria, we conduct two type of experiments.
+Rationality: Since the modeling of convergence criteria involves certain simplification and assumption, we examine the rationality through comparing the actual APR values with the theoretical value of $\sigma_{t}$, as the APR during static phase can be regarded as a reference of $\sigma_{t}$ under ideal conditions.
+We collect a static-movement-static sequence with Livox AVIA.
+As shown in Fig.
+8, the actual APR values during the static phases are very close to $\sigma_{t}$, indicating good consistency between theoretical derivation of $\sigma_{t}$ and actual APR results.
+Furthermore, the APR increases and becomes more volatile during the movement phase.
+Too large an APR reflects a higher risk of insufficient convergence.
+We introduce the convergence criteria to prevent the asymptotic distortion compensation from starting with too high an initial error.
+Validity: We examine the accuracy of AC-LIO at different $\eta$ threshold levels for both AVIA and VLP16, and compare with the benchmark of FAST-LIO2.
+While keeping other parameters the same, we only adjust the threshold of convergence criteria.
+The experimental results are shown in Table III and Fig.
+7.
+  • $\eta\in[0,0.5\sigma_{t}]$: No or rare backpropagation and asymptotic compensation is performed under this condition, reflecting the basic odometry accuracy.
+  • $\eta\in[\sigma_{t},2\sigma_{t}]$: Recommended interval in consideration of other error factors, assuring the backpropagation and asymptotic compensation will be performed with sufficient convergence in previous frame.
+  • $\eta\in[2.5\sigma_{t},3\sigma_{t}]$: Too high a threshold may lead to the backpropagation in case of large initial errors, potentially degrade the overall system accuracy and robustness.
+Figure 8: Rationality evaluation of convergence criteria.
+The orange curve refers the average point-to-plane residual (APR) during a static-movement-static sequence in an indoor scenario.
+The APR values under static phases are very closed to the theoretical value of $\sigma_{t}$, indicating good consistency between theoretical derivation and actual APR results.
+The results show that, in most cases, selecting the threshold $\eta$ within the range of $[\sigma_{t},2\sigma_{t}]$ can achieve higher accuracy.
+As the threshold increases from zero, the odometry accuracy initially improves because more sufficiently converged states are allowed to perform backpropagation and asymptotic compensation.
+However, as the threshold increases further, states with larger initial errors are also allowed to backpropagate, which reduces the effectiveness of the asymptotic compensation and gradually impacts odometry accuracy.
+IV-E Time Efficiency Evaluation
+We calculate the runtime of each algorithm during iteration, with the results shown in Table IV and Fig.
+9.
+And the specific runtime curves can be found in Fig.
+7.
+In terms of average processing time, the time efficiency of our AC-LIO framework ranks second place among the tested frameworks, just following the FAST-LIO2 framework and comparable to the lightweight DLIO framework.
+The maximum time for iteration also satisfies realtime requirement (below 100ms for 10Hz LiDAR input).
+Due to the convergence criteria strategy, our method can avoid incurring extra optimization costs in unreliable or unnecessary conditions, thus achieving the balance between system convergence and high time efficiency.
+TABLE IV: Time Consumption Evaluation ($ms$)
 Ours FAST-LIO2 DLIO SLICT Point-LIO
-Min 4.65 3.15 3.43 21.33 11.95
-Max 74.46 28.48 27.01 345.97 62.36
-Mean 14.62 16.09 9.12 61.97 34.31
-Median 11.71 16.68 8.50 51.49 34.98
-Figure 6: The result of time efficiency evaluation.
+Min 4.12 2.89 3.95 22.61 7.83
+Max 70.69 66.38 37.22 460.24 50.84
+Mean 8.55 7.18 8.57 48.62 24.48
+Median 8.316 6.971 8.24 42.91 23.61
+Figure 9: Time efficiency evaluation.
 The cross line inside box denotes the median process time.
-Our AC-LIO achieves the second highest efficiency, only behind the lightweight DLIO framework.
-And the maximum iteration time also meets the real-time requirements.
-IV-C Time Efficiency Evaluation
-We calculate the runtime of each algorithm during iteration, with the results shown in Table III and Fig.
-6.
-In terms of average and median processing time, the time efficiency of our AC-LIO framework ranks second place among the tested frameworks, just following the lightweight DLIO framework, and the maximum time for iteration also satisfies realtime requirement (below 100ms for 10Hz LiDAR input).
-Due to the convergence criteria strategy, our method can avoid incurring extra optimization costs in unreliable or unnecessary conditions, thus achieving the balance between system convergence and high time efficiency.
-IV-D Convergence Criteria Validity Evaluation
-TABLE IV: Convergence Criteria Validity Evaluation - RMSE of ATE ($m$)
-00 0 $0.5\sigma_{t}$ $[\sigma_{t},2\sigma_{t}]$ $3\sigma_{t}$
-Park_3.2
-0.624 0.610 0.582 0.616
-Subway_3.3
-2.841 2.419 1.670 1.788
-Garden_1006
-1.901 1.709 1.796 1.892
-Garden_1008
-1.197 1.199 1.152 1.178
-To evaluate the validity of the proposed convergence criteria, we examine the odometry accuracy at different $\eta$ threshold levels.
-While keeping other parameters the same, we only adjust the threshold of convergence criteria.
-Due to the factors such as dynamic objects, cumulative pointcloud map errors, etc., we fine-tune the threshold $\eta$ within the interval of $[\sigma_{t},2\sigma_{t}]$.
-The specific experimental results are shown in Table IV.
-  • $\eta=0$: No backpropagation and asymptotic compensation is performed at this situation, reflecting the basic accuracy of the odometry.
-  • $\eta=0.5\sigma_{t}$: Lower than the expected mean of pointcloud residuals, rarely perform backpropagation. (Only considering LiDAR ranging errors, $\mathbb{E}({z})=\sigma_{t}$)
-  • $\eta\in[\sigma_{t},2\sigma_{t}]$: Recommended interval in consideration of other error factors, assuring the backpropagation and asymptotic compensation will be performed with sufficient convergence in previous frame.
-  • $\eta=3\sigma_{t}$: Too high a threshold may lead to backpropagation in case of large initial errors, potentially damaging the system’s accuracy and robustness.
-The experimental results show that, in most cases, selecting the threshold $\eta$ within the range of $[\sigma_{t},2\sigma_{t}]$ achieved higher accuracy.
-Overall, as the threshold increases from zero, the odometry accuracy initially improves because more sufficiently converged states are allowed to perform backpropagation and asymptotic compensation.
-However, as the threshold increases further, states with larger initial errors are also allowed to backpropagate, which reduces the effectiveness of the asymptotic compensation and negatively impacts odometry accuracy.
-V Conclusion and Future Work
-In this paper, we proposed AC-LIO, a novel asymptotically and consistently converging LIO framework.
+Our AC-LIO achieves the second highest efficiency, only behind the FAST-LIO2 and comparable to the lightweight DLIO.
+The maximum runtime also meets the real-time requirements.
+V Conclusion
+In this paper, we proposed AC-LIO, a novel discrete state LIO framework with selective intra-frame smoothing.
 We reconsider the factors that may introduce errors in long-duration, large-scale scenarios, aiming to further enhance the convergence of LIO system.
-The approach of guiding asymptotic compensation based on convergence criteria is designed to promote consistent convergence of pointcloud registration, further improving the odometry accuracy while ensuring system robustness.
-Numerous experiment results indicate that our framework’s average RMSE is reduced by about 30.4% over the second best result, achieving higher accuracy in long-term, large-scale localization and mapping.
-Future Work: We notice the performance of back propagation depends on the guidance of convergence criteria to some extend.
-The introduction of convergence criteria is essentially an attempt to overcome the disruption of the initial error on the back propagation during the iteration update.
-In future work, we will continue to study the convergence assessment of odometry, the combination of asymptotic backpropagation with other strategies, and the perception and adaptation of degenerate scenarios, to further explore the error-analyzable and scenario-interpretable SLAM framework.
+The approach of guiding asymptotic compensation based on convergence criteria is designed to further suppress the motion distortion within LiDAR frame, improving the odometry accuracy while ensuring system robustness.
+Extensive experiments indicate that our framework achieves an average RMSE reduction of approximately 30.4% over the second best method, demonstrating higher accuracy in long-term, large-scale localization and mapping.
+In terms of time efficiency, our framework ranks second place among the evaluated frameworks, just following the FAST-LIO2 and comparable to the lightweight DLIO, achieving a balance between accuracy and real-time performance.